### v1

# Cognitive Paradigms for Evaluating VLMs on Visual Reasoning Task

Abstract
Evaluating the reasoning capabilities of Vision-Language Models (VLMs) in complex visual tasks provides valuable insights into their potential and limitations. In this work, we assess the performance of VLMs on the challenging Bongard Openworld Problems benchmark, which involves reasoning over natural images. We propose and evaluate three human-inspired paradigms: holistic analysis (global context processing), deductive rule learning (explicit rule derivation and application), and componential analysis (structured decomposition of images into components). Our results demonstrate that state-of-the-art models, including GPT-4o and Gemini, not only surpass human benchmarks but also excel in structured reasoning tasks, with componential analysis proving especially effective. However, ablation studies reveal key challenges, such as handling synthetic images, making fine-grained distinctions, and interpreting nuanced contextual information. These insights underscore the need for further advancements in model robustness and generalization, while highlighting the transformative potential of structured reasoning approaches in enhancing VLM capabilities.
Cognitive Paradigms for Evaluating VLMs on Visual Reasoning Task
1 Introduction
Human intelligence seamlessly integrates visual perception with language and higher-level cognition, enabling reasoning about visual information Kunda (2020). This ability, known as visual reasoning, supports tasks such as understanding spatial relationships, inferring object functions, and grounding language in the physical world. Despite advancements in natural language processing (NLP) Zhao et al. (2023), achieving robust visual reasoning in machines remains a significant challenge, particularly in multimodal contexts Cao et al. (2024).
Large Language Models (LLMs) have transformed NLP by excelling in tasks like machine translation, text generation, and question answering Brown et al. (2020); Devlin (2018). Their emergent abilities in areas such as code generation Chen et al. (2021) and logical reasoning Wei et al. (2022) suggest potential for bridging language and vision to tackle complex visual reasoning tasks.
Visual reasoning tasks range from basic object recognition Lin et al. (2014) to abstract inference Zerroug et al. (2022); Teney et al. (2020); Barrett et al. (2018); Zhang et al. (2019). Complex tasks demand integrating perceptual data with higher-order cognition, such as spatial reasoning and relational understanding Cao et al. (2024); Vaishnav et al. (2022). Bongard Problems (BPs) Bongard (1968), involving abstract rules distinguishing two image sets, exemplify such challenges by requiring both perceptual and conceptual reasoning.
Structured reasoning tasks like BPs require the reasoning process to follow specific rules or structures. These tasks involve identifying patterns or relationships among image components and applying abstract reasoning based on those patterns. Vision-Language Models (VLMs), such as Phi-1.5 Javaheripi et al. (2023), Llama-vision Meta (2024), Gemini Google (2024), and GPT-4o OpenAI (2024), provide a robust foundation for tackling BPs. By leveraging multimodal datasets, these models combine visual perception with higher-order cognitive processes, making them well-suited for structured reasoning tasks like BPs. This capability allows them to approach complex visual reasoning by breaking down information into identifiable, logical steps.
Figure 1: Illustration of a Bongard OpenWorld dataset showing three positive and negative samples along with a query image. Distinguishing rule in this example is A group photo at a wedding reception.. Based on the rule, query image belongs to negative category.
This paper investigates the ability of Vision-Language Models (VLMs) to solve natural image-based Bongard Problems using the Bongard Openworld Dataset Wu et al. (2024) (example shown in Figure 1). Inspired by human problem-solving strategies, we evaluate three paradigms: holistic analysis, deductive rule learning, and componential analysis. Our findings demonstrate that VLMs employing these strategies can surpass human performance on this challenging visual reasoning task. Additionally, through ablation studies and a comparative analysis of open- and closed-source models, we identify key limitations in multi-image reasoning, such as difficulties with fine-grained distinctions, synthetic images, and nuanced contextual understanding. These insights provide critical directions for advancing the robustness and human-like reasoning capabilities of VLMs.
Research Contributions:
  1. 1.
    We evaluate the effectiveness of Vision-Language Models (VLMs) on the challenging task of solving Bongard Openworld Dataset—natural image-based BPs.
  2. 2.
    We propose and evaluate three distinct paradigms for applying VLMs to this task, inspired by human learning and problem-solving strategies:
    1. 2.1.
      Holistic Analysis (Analyzing all information together)
    2. $2.$
      Deductive Rule Learning (Focusing on deriving rules before testing scenarios)
    3. 2.3.
      Componential Analysis (Decomposing images into structured, interpretable components)
  3. 3.
    We demonstrate that VLMs can surpass human performance in this challenging open-world visual reasoning task using human-centric learning methodologies.
  4. 4.
    We provide a detailed evaluation of both open-source and closed-source models, highlighting current limitations and performance discrepancies in structured reasoning tasks.
  5. 5.
    We demonstrate that addressing limitations in fine-grained distinctions, synthetic image processing, and contextual understanding is key to improving VLM robustness and aligning them more closely with human reasoning capabilities.
2 Related Work
To evaluate the capabilities of VLMs, various benchmarks have been developed, focusing on different aspects of visual understanding and reasoning. These benchmarks play a crucial role in assessing model performance and identifying areas for improvement. Traditional benchmarks focussed on single image tasks like image captioning Lin et al. (2014), OCR Mishra et al. (2019); Singh et al. (2021), and VQA Antol et al. (2015); Zhang et al. (2024); Lu et al. (2022); Gurari et al. (2018). For multi-image comprehension, researchers have developed interleaved image-text corpora, including MMC4 Zhu et al. (2024) and OBELICS Laurençon et al. (2024); Lin et al. (2024), along with dedicated benchmarks like Q-Bench Wu et al. (2023), Memontos Wang et al. (2024), and DEMON Li et al. (2023). While these benchmarks have contributed to progress in evaluating multi-image understanding, they often lack comprehensive coverage of complex reasoning scenarios. Other benchmarks like MMIU (Multimodal Multi-Image Understanding) Meng et al. (2024) have been pivotal in evaluating VLMs comprehensively, incorporating diverse datasets covering areas like science and mathematics.
Despite advancements in VLM benchmarks that involve language, purely visual reasoning tasks without linguistic input remain crucial for evaluating core visual understanding. Bongard problems Bongard (1968) present a unique challenge in this regard. Their significance lies in their focus on abstract visual concepts, the absence of explicit language, and the inherent multi-image relational reasoning required for their solution. Extensions such as Bongard-HOI Jiang et al. (2022) and Bongard-OpenWorld Wu et al. (2024), further extend these challenges to human-object interactions and real-world scenarios, highlighting models’ few-shot reasoning capabilities.
Beyond Bongard problems, other purely visual reasoning challenges have been proposed, each focusing on different aspects of visual cognition. These include Synthetic Visual Reasoning Test (SVRT) Fleuret et al. (2011), Compositional Visual Reasoning (CVR) Zerroug et al. (2022), Raven’s Progressive Matrices (RPM) Barrett et al. (2018); Zhang et al. (2019), Visual Progressive Matrices (V-PROM) Teney et al. (2020), and the Abstract Reasoning Corpus (ARC) Chollet (2019). These benchmarks collectively assess a range of visual reasoning abilities, including spatial reasoning, compositional reasoning, and abstract pattern recognition, providing valuable insights into the visual understanding capabilities of AI models.
Human cognitive processes offer valuable insights for evaluating visual reasoning in VLMs. Three primary learning strategies are particularly relevant: holistic analysis, deductive analysis, and componential analysis. Holistic analysis involves integrating all available information simultaneously, akin to similarity-based strategies where new information is compared to existing knowledge von Helversen et al. (2014); Johnson-Laird (2010). This approach emphasizes understanding the overall context and interconnections between elements. Focused problem-solving prioritizes understanding underlying principles through deductive reasoning before application, mirroring rule-based strategies Knauff and Wolf (2010); Helie et al. (2015); Rips (1994). This strategy emphasizes abstract rule extraction and application.
Finally, componential analysis breaks down problems into individual components, extracting relevant details before integration Wu and Molnár (2022); Gluck et al. (2008); Hutchinson (1992). This approach emphasizes detailed analysis of individual elements and their subsequent combination. These distinct human learning strategies provide a framework for designing and interpreting visual reasoning benchmarks for VLMs, allowing researchers to assess how well these models emulate different aspects of human cognition.
3 Models
To evaluate our approach comprehensively, we employed a diverse range of Vision-Language Models (VLMs) that include both closed-source and open-source architectures. This diversity allowed us to assess the strengths and limitations of different model types and scales.
This selection encompasses state-of-the-art multimodal models, including GPT-4o OpenAI (2024), Gemini-2.0 Google (2024), Pixtral-12B Agrawal et al. (2024), Llama-Vision-3.2 (11B, and 90B) Meta (2024), LLaVA (7B, 13B and 34B) Liu et al. (2023), and LLaVA built on Llama 3 (8B) XTuner (2025). By leveraging this variety, we systematically analyzed model performance across different architectural designs and parameter scales, offering a holistic view of their capabilities. For the evaluation purpose, two metrics are used: classification accuracy and semantic similarity (more in Appendix C).
Further technical details, including model architectures and individual summaries, are elaborated in Appendix D.
4 Dataset
We evaluate our approach on the Bongard OpenWorld dataset Wu et al. (2024), a benchmark designed to assess few-shot visual reasoning abilities. The full dataset contains 1001 cases, from which we have chosen a sample of 500 cases for our experiments. Each case in the dataset consists of six positive and six negative examples and a query image. These cases are governed by an underlying "commonsense" rule, with the rules themselves categorized into distinct visual concepts (further details can be found in Appendix B). Figure 1 illustrates the task: given a set of positive and negative examples, the goal is to determine the underlying rule and classify a query image accordingly.
5 Experiments
This paper investigates VLMs’ ability to solve natural image-based Bongard Problems. We evaluate three paradigms: holistic analysis (processing all information together), deductive rule learning (deriving rules first), and componential analysis (examining components separately).
5.1 Holistic Analysis
In this section, we describe our experiments using a holistic analysis approach, where the model is presented with all images (positive, negative, and test) simultaneously.
The holistic analysis experiments assess the ability of VLMs to directly infer patterns and rules from a set of images. We designed a prompt that presents the model with m positive examples, n negative examples, and a final query image.
The prompt instructs the model to:
  1. 1.
    Identify a rule that distinguishes the positive examples from the negative examples;
  2. 2.
    Analyze the query image;
  3. 3.
    Classify the query image as either positive or negative based on the identified rule.
It is explicitly requested to the VLM model to provide its analysis, the derived rule, a description of the query image, and the final classification category. Details are provided in Appendix E.1.
5.2 Deductive Rule Learning
This section describes our experiments using a deductive rule learning approach, which involves a two-step process: first, the model infers a general rule; second, it applies this rule to classify a query image.
The deductive rule learning experiments assess the ability of VLMs to explicitly formulate a rule before classifying a query image. This approach mirrors human deductive reasoning. We employed a two-stage prompting strategy. In the first stage, we presented the model with m positive examples and n negative examples and asked it to identify a rule that distinguishes the two categories. The prompt requested a concise summary of the identified rule (less than 20 words). The precise wording of this first-stage prompt is provided in Appendix E.2.
In the second stage, we presented the model with the previously generated rule summary and a query image. The prompt instructed the model to analyze the query image in the context of the given rule and classify it as either positive or negative. The prompt also requested a brief analysis of the query image and a restatement of the rule. The precise wording of this second-stage prompt is provided in Appendix E.2.
5.3 Componential Analysis
This section describes our experiments using a componential analysis approach (detailed in Appendix E.3). This method involves decomposing each image into its constituent visual elements and then reasoning based on these components (textual elements). This process is divided into three main stages: (1) detailed image description generation, (2) providing instruction on rule derivation, and (3) rule extraction and application based on these descriptions.
The componential analysis experiments aim to evaluate the ability of VLMs to reason based on structured descriptions of visual content. This approach simulates how humans might break down complex visual scenes into simpler parts for analysis. We employed a three-stage prompting strategy.
In the first stage, we generate detailed descriptions of each image. This prompt instructs the model to analyze an image and provide a structured description, covering aspects such as scene, objects (living and inanimate), activities, contextual elements, visual patterns, emotional undertones, textual information, and a summary. In the second stage, we provide the models with instructions on how to analyze the outputs and derive the rule.
In the final stage, which is purely textual, we provided the model with the descriptions generated in the first stage, along with information about which images belong to positive examples and negative examples, and the query image. The prompt then asks the model to derive a rule based on the descriptions of positive and apply this rule to classify the query image.
5.4 Rule-Based Evaluation
This section describes experiments designed to disentangle the image analysis (perception) and rule application (abstract reasoning) capabilities of VLMs. Unlike the previous experiments, where the models were tasked with both extracting and applying rules, here we provide the models with pre-extracted rules and assess their ability to apply these rules to classify query images.
These experiments isolate the rule application component of the Bongard problem-solving process. This experimental setup is conceptually similar to the Deductive Rule Learning experiments (Section 5.2), which also involved a two-step process: (1) inferring a general rule and (2) applying that rule to classify a query image. However, in this experiment, we bypass the first step (rule inference) and instead use the rule summaries generated by GPT-4o in the Deductive Rule Learning experiment. For each Bongard problem, we took the rule summary generated by GPT-4o. We then presented the model with the query image and the corresponding rule summary, explicitly instructing it to classify the query image based on the provided rule. This approach allows us to specifically assess how well the models can understand and apply a given rule, independent of their ability to extract the rule themselves.
6 Results
6.1 Holistic Analysis
Holistic analysis necessitates models with multi-image and multimodal capabilities. Among the models tested, only Gemini 2.0 (flash) (referred to Gemini here) and GPT-4o possess these features.
Table 1 presents the overall accuracy, as well as the accuracy on positive and negative samples separately. As shown, Gemini achieves a slightly higher overall accuracy of 82.2% compared to GPT-4o’s 80.0%. Unlike GPT-4o, Gemini shows a more balanced performance across both positive and negative samples, suggesting a more robust approach to handling cases where no clear rule might exist for the negative examples.
These results suggest that while GPT-4o may excel at recognizing patterns within the positive examples, its lower accuracy on negative samples indicates challenges in distinguishing subtle counter examples. On the other hand, Gemini’s balanced performance reflects greater robustness in addressing the diverse reasoning demands of holistic analysis tasks.
Model Holistic Analysis Deductive Rule Learning Componential Analysis
neg pos Overall neg pos Overall neg pos Overall
GPT-4o 66.4 93.6 80.0 82.8 93.2 88.0 92.8 92.8 92.8
Gemini 2.0 80.48 83.94 82.2 85.66 87.95 86.8 90.84 96.39 93.6
Pixtral-12B - - - - - - 59.2 48.4 53.8
Llava-Llama3-8B - - - - - - 93.02 13.08 53.15
Llama-vision-11B - - - - - - 42.8 64.0 53.4
Llava 7B - - - - - - 65.20 67.20 66.20
Human - - - - - - - - 91.0
Table 1: Classification accuracy across different Analysis Paradigms and Human performance Wu et al. (2024). neg and pos represents the classification accuracy for negative and positive query images.
6.2 Deductive Rule Learning
This two-stage approach allows us to evaluate the model’s ability to not only identify patterns but also articulate them as explicit rules, providing insights into its deductive reasoning capabilities. We conducted these experiments using Gemini and GPT-4o. Table 1 summarizes the overall accuracy, as well as the accuracy on positive and negative samples separately.
This time GPT-4o achieves a slightly higher overall accuracy compared to Gemini. Both models demonstrate balanced performance across positive and negative samples. GPT-4o, in particular, excels in classifying positive samples, while Gemini achieves a more consistent performance across both categories.
These results indicate that the two-stage deductive approach enhances the models’ ability to handle the complexities of both positive and negative examples, surpassing the performance seen in the holistic approach. GPT-4o’s superior accuracy on positive samples suggests its effectiveness in identifying and applying explicit patterns, while Gemini’s balanced results highlight its robustness across diverse reasoning scenarios.
6.3 Componential Analysis
The componential analysis approach provides a framework for evaluating a broader spectrum of vision-language models (VLMs) by decomposing the complex multi-image reasoning task into manageable single-image description tasks. This design facilitates the inclusion of models that lack explicit multi-image processing capabilities but can effectively analyze individual images and process structured textual representations.
We conducted the three-stage componential analysis experiment using six VLMs: Pixtral-12B, Llava-Llama3-8B, Llama-Vision-11B, Llava 7B, GPT-4o, and Gemini. The performance metrics, presented in Table 1, include overall accuracy and accuracy on positive and negative samples.
The results demonstrate a wide range of performance across the evaluated models. GPT-4o and Gemini achieve near-optimal accuracies, with overall accuracies of 92.8% and 93.6%, respectively. Notably, these results surpass human-level performance, which is reported at 91% on this task. This finding highlights the potential of advanced VLMs to excel in tasks traditionally dominated by human reasoning capabilities.
Open-source models, including Pixtral-12B, Llava-Llama3-8B, and Llama-Vision-11B, Llava 7B exhibit significant variability. Llava-Llama3-8B, for example, achieves a high accuracy on negative samples but an exceptionally low accuracy on positive samples, indicating a strong bias toward negative classifications. Similarly, Llama-Vision-11B demonstrates an inverse pattern. These discrepancies suggest that these models struggle to generalize effectively across different classes.
The superior performance of GPT-4o and Gemini indicates that structured, detailed descriptions significantly enhance reasoning capabilities in VLMs. Their ability to surpass human-level performance underscores the importance of integrating multimodal capabilities with robust processing pipelines. Furthermore, performance open-source models highlights the potential for continued advancements in accessible VLMs to achieve substantial progress in complex reasoning tasks.
6.4 Rule-Based Evaluation
The Rule-Based Evaluation experiment isolates the models’ ability to apply explicitly provided rules to classify query images, bypassing the rule inference step. This approach offers a focused assessment of the rule application capabilities of the models. The results, presented in Table 2, reveal distinct trends across the evaluated models.
Pixtral-12B achieved the highest overall accuracy (83.8%), indicating a strong ability to comprehend and operationalize the provided rules effectively. Its relatively balanced performance on both negative and positive samples highlights its robust application of rules across varying contexts.
Model Accuracy (%)
Overall neg pos
LLaVA-7B 66.20 65.20 67.20
Llama-vision-11B 68.2 38.0 98.4
Pixtral-12B 83.8 75.2 92.4
LLaVA-13B 70.0 58.8 81.2
LLaVA-34B 74.8 52.8 96.8
Llama-vision-90B 74.2 59.6 88.8
Table 2: Rule-Based Evaluation Results showing overall classification accuracy along with categorical accuracy for positive (pos) and negative (neg) samples.
The results in Table 2 highlight that while certain models excel in confirming positive matches, their difficulties in handling negative cases expose potential biases in rule application mechanisms. Pixtral-12B’s relatively balanced performance suggests it is more adept at applying rules across diverse scenarios, while models like Llama-Vision-xx and LLaVA-xx may overfit to confirmatory evidence.
By isolating the rule application component, this experiment provides valuable insights into the models’ strengths and limitations. When combined with findings from the Deductive Rule Learning experiments (Section 5.2), this analysis helps to disentangle the factors contributing to model performance, distinguishing between deficits in rule extraction and rule application. Such insights are crucial for guiding improvements in visual-language model reasoning capabilities.
7 Ablation
7.1 Semantic Similarity Analysis
To further analyze the reasoning process of the best-performing models, GPT-4o and Gemini, we conducted a semantic similarity analysis.This analysis aims to quantify the alignment between the descriptions of the query images and the rules identified for each category. We hypothesized that for correctly classified positive examples, the similarity between the description and rule should be high, while for negative examples, it should be lower.
[…]
7.2 Analysis per Commonsense Category
To better understand the comparative capabilities of GPT-4o and Gemini, we analyzed their performance across ten commonsense reasoning categories in the Bongard Openworld dataset (details in Appendix B.3). This analysis (in Table A.5) highlights the models’ strengths and limitations in handling diverse commonsense reasoning categories. Both models demonstrated robust performance across most categories indicating models’ capability to process high-level conceptual patterns effectively.
Differences in performance were observed in categories requiring more nuanced or contextual understanding, such as "Taste / Nutrition / Food" and "Human-Object Interaction." This suggests that while both models are adept at general reasoning tasks, their architecture and training data influence their performance on specific types of visual information. Notably, Gemini showed strengths in scenarios involving fine-grained object attributes, while GPT-4o performed better in tasks demanding contextual interpretation.
7.3 Analysis of Incorrectly Classified Samples
In this ablation study, we analyzed the misclassified samples from GPT-4o and Gemini during the Componential Analysis experiments to identify common limitations. The total 15 common test samples were incorrectly identified by both the VLMs. The errors revealed several recurring issues (more details in Table A.7), including challenges with synthetic images that lack natural textures, difficulties in making fine-grained distinctions, and struggles with recognizing specific image types, such as satellite views. Additionally, the models showed limitations in understanding nuanced contextual information, as evidenced by misclassifications related to artistic styles and the broader scene context. These findings highlight that while the models excel in many areas, they still face challenges in handling complex visual details and contextual reasoning, suggesting areas for improvement in VLMs’ robustness and generalization.
8 Discussion
Our study investigates the application of Vision-Language Models (VLMs) to Bongard Open-World Problems (BPs), providing insights into the computational mechanisms underlying visual reasoning and their alignment with human cognitive strategies. We introduced three evaluation paradigms—Holistic Analysis, Deductive Rule Learning, and Componential Analysis—designed to mirror distinct human approaches to learning and problem-solving. These paradigms allow us to evaluate VLM performance and understand how these models relate to human cognition.
A critical aspect of human scene understanding is the dynamic interaction between the observer and their environment Zelinsky (2013); Malcolm et al. (2016); Gibson (2015). This interactionist view posits that understanding a scene is an active process shaped by behavioral goals, rather than a passive reception of sensory input. This active engagement with the visual world is directly relevant to how both humans and VLMs approach BPs. Rather than processing individual images in isolation, both humans and VLMs benefit from considering the relationships between images in a given set.
Our Holistic Analysis paradigm specifically targets this relational aspect by requiring VLMs to process all images and descriptions simultaneously. This approach allows the model to capture the overall context and directly identify the distinguishing rule, akin to human holistic problem-solving, where the "gist" of a scene is rapidly perceived, guiding subsequent attention and actions Biederman (1987); Li et al. (2002). Such holistic processing provides a crucial initial understanding, which can then be refined through more detailed analysis. This process heavily relies on working memory to maintain the "gist" while further analysis unfolds Baddeley (2012).
Following this initial holistic understanding, human visual information processing is further refined through attentional mechanisms that guide gaze and cognitive processing Posner (1980). While eye movements can reflect attentional shifts, our Deductive Rule Learning paradigm directs the VLMs’ attention on learning the underlying rule prior to evaluating query images. This mirrors human deductive reasoning, where a hypothesis is formed and then tested against new evidence. By focusing on rule inference before rule application, the VLM can store learned rules in long-term memory Squire (1992), allowing for efficient application to unseen instances.
In addition to holistic and deductive strategies, humans often use analytical approaches to break down complex problems. The Componential Analysis paradigm reflects this tendency by decomposing a complex problem into smaller, more manageable components. In this paradigm, VLMs analyze positive and negative examples separately, extracting individual rules for each set before comparing them to identify the distinguishing factor. This process aligns with cognitive theories that distinguish between System 1 (fast, intuitive perception) and System 2 (slow, deliberate reasoning) Kahneman (2011). Object recognition and categorization, a System 1 process involving perceptual mechanisms Marr (2010), precedes the more deliberate System 2 reasoning required to compare and contrast extracted rules. This analytical decomposition facilitates a more precise understanding of the underlying rules.
Our findings, which show that VLMs can surpass human performance on BPs when applying these human-inspired paradigms, underscore the importance of attention and memory mechanisms in visual reasoning. Attention mechanisms, critical for selecting relevant information, are implemented in VLMs through architectures such as those described in Bahdanau et al. (2016) and Vaswani et al. (2017). These mechanisms allow VLMs to selectively focus on relevant image features, similar to human gaze patterns and attentional shifts. Meanwhile, memory mechanisms are essential for encoding visual information, maintaining it during rule formulation, and retrieving it for comparison. This is evident in the VLMs’ ability to store learned rules (long-term memory) and maintain representations of images during analysis (working memory). The interplay between attention and memory is crucial for both human and artificial intelligence, providing deeper insights into the computational foundations of visual reasoning Vaishnav and Serre (2023).
Moreover, our evaluation of both open-source and closed-source models highlights significant performance discrepancies, shedding light on the limitations of current VLMs in handling complex visual reasoning tasks. These limitations underscore the need for further model refinement, particularly in handling multi-image reasoning tasks and improving rule extraction capabilities. Finally, we demonstrate that effective reasoning in complex visual scenarios requires the complementary strengths of VLMs and Large Language Models (LLMs). While VLMs excel at extracting features from images, LLMs play a crucial role in higher-level reasoning tasks, indicating a promising direction for future research where the integration of both types of models can enable more sophisticated problem-solving in visual reasoning tasks.
9 Limitation
Despite the promising results of this study, several limitations must be considered. The Bongard Openworld dataset may not fully capture the complexity and diversity of real-world visual reasoning tasks, potentially limiting generalizability. Additionally, the reliance on pre-extracted rule summaries in some experiments restricts the evaluation of models’ complete reasoning capabilities, particularly in rule extraction. The scalability of large models, such as GPT-4o and Gemini, is also a concern due to their high computational demands, and their ability to generalize to noisy, real-world data remains uncertain. Furthermore, the opacity of model decision-making processes highlights the need for better interpretability, and the risk of overfitting to task-specific knowledge could hinder broader applicability. Lastly, our evaluation primarily focuses on accuracy and semantic similarity, which may not fully capture the nuances of model reasoning. These limitations highlight areas for future research and refinement in applying VLMs to complex reasoning tasks.
10 Conclusion
This work investigates the effectiveness of Vision-Language Models (VLMs) in solving the challenging Bongard Openworld dataset, a task requiring complex visual reasoning. We proposed three human-inspired paradigms—Holistic Analysis, Deductive Rule Learning, and Componential Analysis—and demonstrated that VLMs, particularly GPT-4o and Gemini, can surpass human-level performance. Our results highlight that VLMs excel when combining detailed image analysis with structured rule-based reasoning. The study also emphasizes the complementary roles of VLMs and Large Language Models (LLMs), with VLMs excelling at visual feature extraction and LLMs crucial for higher-level reasoning. These findings provide valuable insights into model design and evaluation, suggesting a promising direction for future research in complex visual reasoning tasks and real-world applications.
Acknowledgments
This work was supported by the Estonian Centre of Excellence in AI [grant number TK213U8]
Appendix
Appendix A Broader Relevance
Broader Relevance
The insights gained from this work have broader implications beyond the specific task of solving the Bongard Openworld dataset. As the capabilities of Vision-Language Models (VLMs) continue to evolve, the methodologies and paradigms explored in this paper can contribute significantly to the development of models capable of more advanced reasoning and problem-solving in real-world applications. The ability to tackle complex visual reasoning tasks, such as the Bongard problems, is critical for a range of fields, from autonomous systems and robotics to medical imaging and beyond.
One of the most significant implications of our work is the demonstration that VLMs, when integrated with structured reasoning paradigms, can outperform human-level performance in specific problem-solving scenarios. This opens up new avenues for the deployment of AI systems in environments that require a blend of visual perception and reasoning, such as dynamic real-world settings where an agent must continually adapt to new and diverse scenarios. By surpassing human-level performance in tasks that require complex visual analysis, VLMs could enable more robust and adaptive AI systems that can be deployed in areas ranging from automated decision-making to real-time problem solving.
Furthermore, our work also highlights the critical gap between the reasoning abilities of VLMs vs Large Language Models (LLMs). A combination of both is essential is needed for complex reasoning tasks. While VLMs excel at understanding and processing visual inputs, LLMs bring the reasoning and linguistic abilities necessary to contextualize and apply rules based on these inputs. This integration is essential for advancing research in cognitive modeling, where understanding the interplay between perception, reasoning, and language is central. Our findings encourage further exploration of hybrid architectures that seamlessly blend visual and language models, potentially driving progress in areas such as visual question answering, image captioning, and multimodal learning.
Finally, our detailed evaluation of both open-source and closed-source models contributes to the ongoing dialogue about the accessibility, fairness, and transparency of AI systems. By examining the performance discrepancies across various architectures, we not only illuminate the current limitations of VLMs but also provide actionable insights for future improvements. This work could inspire the development of more efficient and capable models, particularly in the open-source community, and contribute to efforts aimed at making advanced AI more accessible and usable for a wide range of applications.
In summary, this work sets the stage for more advanced, human-like AI systems capable of solving complex visual reasoning tasks, with far-reaching applications in domains such as autonomous systems, healthcare, and cognitive modeling. The methodologies and findings outlined here provide a valuable foundation for future research and development in AI, pushing the boundaries of what is possible with multimodal learning and reasoning.
Appendix B Dataset
B.1 Bongard OpenWorld Dataset
We utilize a subset of 500 cases from the Bongard OpenWorld dataset Wu et al. (2024), a benchmark for evaluating few-shot reasoning in machine vision. Each case presents a set of positive and negative examples, distinguished by an underlying "commonsense" rule. These rules, represented by numerical values (0-9), encode different visual and spatial concepts, as detailed below.
B.2 Commonsense Value Categories
The Bongard OpenWorld dataset categorizes the underlying rules into several visual concept categories, some of which are explicitly related to commonsense reasoning. Table A.3 provides a summary of these categories and example concepts.
Concept Category ID Example
Anything else 0 Animals are running.
Human-Object Interaction (HOI) 1 A person playing the guitar.
Taste / Nutrition / Food 2 A plate of high-calorie food.
Color / Material / Shape 3 A wooden floor in the living room.
Functionality / Status / Affordance 4 An animal capable of flying in the tree.
And / Or / Not 5 A man without beard.
Factual Knowledge 6 A building in US capital.
Meta Class 7 Felidae animals.
Relationship 8 A bench near trees.
Unusual Observations 9 Refraction of light on a glass cup.
Table A.3: Commonsense ID Categories in Bongard OpenWorld
B.3 Commonsense Value Distribution in Our Subset
Table A.5 provides a detailed breakdown of the distribution of commonsense values within our 500 test-cases subset. As observed, the value ’0’ ("Anything else," marked as commonsense) is significantly more prevalent than other values. This distribution should be considered when analyzing the model’s performance, as it may indicate a bias towards learning certain types of rules, particularly those related to the "Anything else" category.
ID Count Percentage
0 365 73%
3 36 7.2%
9 26 5.2%
1 15 3%
2 14 2.8%
4 12 2.4%
5 10 2%
6 10 2%
8 8 1.6%
7 4 0.8%
Total 500 100%
Table A.4: Distribution of Commonsense ID in Bongard OpenWorld Dataset
B.4 Dataset Availability
The complete Bongard OpenWorld dataset is publicly available at Bongard-OpenWorld.github.io.
Appendix C Evaluation Metrics
Classification Metric:
For evaluating the performance of the VLMs, we employ categorization accuracy on the binary classification task. This metric measures the percentage of test instances correctly classified as either positive or negative as present in the Bongard Problems.
Semantic Similarity:
We leverage a Semantic Similarity Evaluator to quantify the semantic relatedness between generated responses (e.g., question answers) and ground truth answers. This evaluator is inspired by the work of Risch et al. (2021), which explores semantic answer similarity for evaluating question answering models.
Appendix D Model Details
This section provides comprehensive details on the Vision-Language Models (VLMs) utilized in our experiments.
D.1 Vision-Language Models (VLMs)
  • ChatGPT (GPT-4o with vision): OpenAI’s GPT-4o, accessed with vision capabilities via the openai library, processes both images and text seamlessly. It OpenAI (2024) excels in reasoning and instruction-following tasks, offering robust text generation and broad general knowledge. However, its closed-source nature and cost make it less accessible for extensive experimentation.
  • Gemini (2.0-flash-exp): Google’s Gemini is a multimodal model Google (2024) emphasizing speed and computational efficiency in its "flash" version, accessed through the google.generativeai library. While it provides rapid inference capabilities, it is an experimental model with limited public documentation, restricting detailed architectural insights.
  • Pixtral (12B): Accessed via the mistralai library, Pixtral is a 12-billion-parameter model designed for vision-language tasks Agrawal et al. (2024). Its architecture and training details are not fully disclosed, but it shows strong performance in image-text reasoning. As a proprietary model, its access is constrained to specific APIs.
  • Llama-Vision (3.2, 11B, and 90B): Llama-Vision Meta (2024) combines a vision encoder with the Llama language model, with experiments conducted on 11B and 90B parameter variants to evaluate the impact of scaling. Accessed via the ollama library, its open-source availability enables detailed exploration, but the 90B version requires significant computational resources.
  • LLaVA (7B, 13B and 34B): LLaVA (Large Language and Vision Assistant) integrates a vision encoder with a large language model, fine-tuned for visual instruction-following tasks Liu et al. (2023). Accessed through the ollama library, it is particularly effective in visual question answering but demands substantial GPU memory.
  • LLaVA-Llama 3 (8B): This variant of LLaVA incorporates the Llama 3 architecture XTuner (2025). Accessed via the ollama library, it benefits from the advancements of Llama 3 Dubey et al. (2024), offering improved efficiency over larger models. However, its relative novelty means limited community feedback and tooling support.
D.2 Model Configurations and Hyperparameters
For closed-source models, we utilized API calls by installing the openai, mistralai (for Pixtral), and google.generativeai libraries, enabling access to ChatGPT (GPT-4o with vision) OpenAI (2024), Pixtral (12B) Agrawal et al. (2024), and Gemini (2.0-flash-exp) Google (2024), respectively. The open-source models, on the other hand, were accessed using the ollama Ollama (2025) library.
  • Input Format: Bongard problems were presented by encoding images in base64 and embedding them within text prompts. For instance: "Here are two sets of images. The first set represents positive examples, and the second set represents negative examples of a concept. What is the rule that distinguishes the two sets? "
  • Image Size: For OpenAI and Gemini models, the default image sizes provided by their APIs were used. For all other models, the maximum image size was restricted to 1024 pixels to ensure uniformity.
  • Decoding Parameters: The temperature parameter was set to 0 to maintain deterministic outputs during evaluation.
  • Fine-tuning: No fine-tuning was performed; all models were evaluated in their pre-trained configurations.
  • Hardware: Experiments were conducted using NVIDIA GPUs, including 2080Ti, 3090, and 6000 Ada, ensuring adequate computational power for all model evaluations.
Appendix E Results
Concept Category ID GPT 4.0 Gemini 2.0
Anything else 0 92.88 94.23
Human-Object Interaction (HOI) 1 86.67 92.86
Taste / Nutrition / Food 2 100.00 85.71
Color / Material / Shape 3 88.89 91.67
Functionality / Status / Affordance 4 100.00 100.00
And / Or / Not 5 90.00 80.00
Factual Knowledge 6 90.00 90.00
Meta Class 7 100.00 100.00
Relationship 8 100.00 100.00
Unusual Observations 9 92.31 92.31
Table A.5: Accuracy per Commonsense ID and Categories in Bongard OpenWorld
Gemini 2.0 GPT-4o
Category Mean Std Mean Std
Positive 0.915 0.02 0.9022 0.02
Negative 0.868 0.02 0.8659 0.02
Table A.6: Semantic Similarity between query image Descriptions and Identified Rules
SL NO Task ID Caption Reason
1 0021_neg_0 Cars on the city streets at night The image is synthetic, not a real-world photograph.
2 0014_neg_0 A person playing a guitar. The image is synthetic, lacking real-world context.
3 0033_neg_0 A bicycle is placed in the corner The image is synthetic and does not represent real-world scenarios.
4 0037_neg_0 The girl has long and thin braids on her head. While technically correct, the braids are incorrectly interpreted or misrepresented.
5 0076_pos_0 Various kinds of rings The classification error arises due to confusion between antique and artistic presentation styles.
6 0076_neg_0 Various kinds of rings The classification error arises due to confusion between antique and artistic presentation styles.
7 0082_neg_0 Live coral on the sea floor. The analysis is too fine-grained.
8 0084_neg_0 A wooden fence surrounding a grassy field. Failure to accurately identify the wooden fence in the image.
9 0112_neg_0 A wooden floor in the living room. The analysis is too fine-grained, and the image does not match the caption.
10 0117_neg_0 Colorful ribbons. While ribbons are present, they are not colorful as described.
11 0122_neg_0 A satellite view of Earth. Unable to recognize or correctly interpret the satellite image.
12 0136_pos_0 Spectator seats view in the stadium. The analysis is overly detailed, leading to a categorization error.
13 0213_neg_0 Checkerboard pattern fabrics Only the checkerboard pattern was identified, missing other key aspects.
14 0234_neg_0 A beautiful stone sculpture The image is misclassified as a stone obelisk monument rather than a sculpture.
15 0247_pos_0 Small river filled with reeds Although the identification is correct, the error stems from overly fine-grained analysis.
Table A.7: Task IDs and descriptions where GPT-4o and Gemini both incorrectly categorize.
E.1 Holistic Analysis
def visual_concept_test_prompt(m, n):
Generates a visual analysis prompt.
Args:
m (int): Number of positive samples.
n (int): Number of negative samples.
Returns:
str: The formatted prompt string.
return f"""
You are provided with {m + n + 1} images: the first {m} samples are ‘cat_2‘, the next {n} samples are ‘cat_1‘, and the last image is the ‘query image‘.
Analyze the common characteristics or patterns found in the ‘cat_2‘ samples (positive samples: following 1 common rule) that distinctly separate them from the ‘cat_1‘ samples (negative samples: it might not follow any possible rule).
Your task is to:
1. Determine the rule or criterion that distinguishes the ‘cat_2‘ samples from the ‘cat_1‘ ones.
2. Analyse the ‘query image‘ (last image).
3. Provide your conclusion for the ‘query image‘ if it can be categorized as either ‘cat_1‘ or ‘cat_2‘ based on the analysis and the rule.
Ensure that the output is clear, well-formatted, and free of unnecessary explanations.
Omit the ‘‘‘ tags at the beginning and end of the page. The format of your output should be as follows:
- **Analysis**: (Your analysis here)
- **Rule**: (The distinguishing rule here)
- **Query Image**: (Query image details)
- **Conclusion**: (cat_1 or cat_2)
E.2 Deductive Rule Learning
First-Stage Prompt (Rule Extraction)
def visual_concept_prompt(m, n):
try:
if m < 0 or n < 0:
raise ValueError(f"Invalid input: m and n must be non-negative. Received m={m}, n={n}.")
if m > 0 and n > 0:
prompt = f"""
You are provided with {m + n} images: the first {m} samples are cat_2, the next {n} samples are cat_1. Analyze the common characteristics or patterns found in the cat_2 samples (positive samples: following 1 common rule) that distinctly separate them from the cat_1 samples (negative samples: it might not follow any possible rule).
Your task is to provide the rules that defines cat_2 samples. At the end, write "summary" of the rule identified in less than 20 words.
Ensure that the output is clear, well-formatted, and free of unnecessary explanations. Omit the ‘‘‘ tags at the beginning and end of the page.
if n == 0:
prompt = f"""
You are provided with {m} images: {m} samples are cat_2. Analyze the common characteristics or patterns found in the cat_2 samples (positive samples: following 1 common rule) that distinctly separate them from negative samples which might not follow any possible rule.
Your task is to provide the rules that defines cat_2 samples. At the end, write "summary" of the rule identified in less than 20 words.
Ensure that the output is clear, well-formatted, and free of unnecessary explanations. Omit the ‘‘‘ tags at the beginning and end of the page.
return prompt
except ValueError as e:
print(f"Error: {e}")
raise
Second-Stage Prompt (Rule Application)
# Define the visual analysis prompt
def visual_concept_test_prompt(m, n, summary):
return f"""
[…]
Your task is to ponder over the rule and provide your conclusion for the ‘query image‘ if it can be categorized as either "cat_1" or "cat_2".
Ensure that the output is clear, well-formatted, and free of unnecessary explanations.
Omit the ‘‘‘ tags at the beginning and end of the page. The format of your output should be as follows:
- **Analysis**: (Your analysis here)
- **Rule**: (The distinguishing rule here)
- **Query Image**: (Query image details)
- **Conclusion**: (cat_1 or cat_2)
E.3 Componential Analysis
First-Stage Prompt (Image Description Generation)
# Define the visual analysis prompt
def visual_concept_prompt():
Generates a visual analysis prompt.
Args:
Returns:
str: The formatted prompt string.
return """
Carefully examine the provided image and identify all possible visual elements, organizing them into a detailed hierarchical structure. Start with broad categories and progress to more specific subcategories. This should cover everything visible in the image, ensuring no detail is overlooked. Structure your findings in a JSON format to enable easy comparison and synthesis of data from other images. This will help discern patterns, contexts, and rules valuable for identifying or understanding query images.
Your hierarchy might encompass the following elements:
1. **Scene/Environment**: Description of the overall setting depicted, such as urban, natural, indoor, or outdoor scenes.
2. **Objects**: Define distinct items or entities present in the scene.
- **Living Beings**: Animals, humans, or other biological entities.
- Species or classification (e.g., dog, bird, human).
- Characteristics (e.g., color, posture, movement).
- **Inanimate Objects**: Both synthetic and natural elements.
- Categories (e.g., vehicle, building, trees).
- Properties (e.g., color, size, material, shape).
3. **Activities**: Observable actions or interactions involving any objects or beings.
- Specific descriptions of actions (e.g., walking, flying).
- Participants involved in these actions.
4. **Contextual Elements**: Environmental conditions and time markers, such as time of day or weather.
- Detailed characteristics (e.g., cloudy, night, winter).
5. **Visual Patterns**: Prominent colors, textures, and patterns that are visually significant.
6. **Emotional Undertones**: Any emotional presence or expressions evident in the image.
7. **Textual Information**: Any visible text within the image, including what it says and its visual style.
8. **Summary**: A concise narrative summarizing the overall content and context of the image.
Ensure that every aspect from the image is represented under these categories. The information should be presented in the following JSON format:
"Scene": {
"Description": "..."
"Objects": {
"Living Beings": [...],
"Inanimate Objects": [...]
"Activities": [...],
"Contextual Elements": {
"Time of Day": "...",
"Weather": "..."
"Visual Patterns": {
"Dominant Colors": [...],
"Textures": [...]
"Emotional Undertones": "..."
"Textual Information": "..."
"Summary": "..."
Ensure that the JSON output is clear, well-formatted, and free of unnecessary explanations. Omit the ‘‘‘json tags at the beginning and end of the page.
Second-Stage Prompt (Rule Derivation Instruction)
def system_eval_prompt():
return """
You are an expert in visual reasoning and analysis, working with complex datasets such as the Bongard dataset. Your task is to analyze descriptions of images from two categories (cat_2 and cat_1) along with a query image, provided in JSON format.
The goal is to:
1. Identify the **common characteristics or patterns** in the positive samples (cat_2) that distinctly separate them from the negative samples (cat_1).
2. Derive a **clear and concise rule** that defines the positive samples.
3. Apply this rule to the query image and categorize it as belonging to either cat_1 or cat_2.
Ensure the output is clear, concise, and follows the provided format:
- **Analysis**: Provide a brief analysis of the distinguishing characteristics or patterns.
- **Rule**: State the rule that separates cat_2 from cat_1.
- **Query Image**: Analyze the query image based on the derived rule.
- **Conclusion**: Clearly state whether the query image belongs to cat_1 or cat_2.
Respond in the format provided, and avoid unnecessary explanations or text.
Third-Stage Prompt (Rule Application and Classification)
def user_eval_prompt(all_image_specs, m, n):
return f"""
We are working with the Bongard dataset, which contains {m} images in cat_2 (positive samples) and {n} images in cat_1 (negative samples). These categories are defined as follows:
- Cat_2: Positive samples that follow a single common rule.
- Cat_1: Negative samples that may not follow any specific rule.
The image descriptions for the positive samples, negative samples, and the query image are provided in JSON format. Analyze the common patterns or characteristics in the cat_2 samples that distinguish them from cat_1 samples.
Your task is to:
1. Derive the rule that defines the cat_2 samples.
2. Apply this rule to categorize the query image.
Here are the image descriptions:
### Positive Samples (cat_2):
{all_image_specs[:m]}
### Negative Samples (cat_1):
{all_image_specs[m:m+n]}
### Query Image:
{all_image_specs[-1]}
Provide your output in the following format:
- **Analysis**: (Your analysis here)
- **Rule**: (The distinguishing rule here)
- **Query Image**: (Query image details)
- **Conclusion**: (cat_1 or cat_2)

### v2

@@
 Abstract
-Evaluating the reasoning capabilities of Vision-Language Models (VLMs) in complex visual tasks provides valuable insights into their potential and limitations.
-In this work, we assess the performance of VLMs on the challenging Bongard Openworld Problems benchmark, which involves reasoning over natural images.
-We propose and evaluate three human-inspired paradigms: holistic analysis (global context processing), deductive rule learning (explicit rule derivation and application), and componential analysis (structured decomposition of images into components).
-Our results demonstrate that state-of-the-art models, including GPT-4o and Gemini, not only surpass human benchmarks but also excel in structured reasoning tasks, with componential analysis proving especially effective.
-However, ablation studies reveal key challenges, such as handling synthetic images, making fine-grained distinctions, and interpreting nuanced contextual information.
-These insights underscore the need for further advancements in model robustness and generalization, while highlighting the transformative potential of structured reasoning approaches in enhancing VLM capabilities.
+Advancing machine visual reasoning requires a deeper understanding of how Vision-Language Models (VLMs) process and interpret complex visual patterns.
+This work introduces a novel, cognitively-inspired evaluation framework to systematically analyze VLM reasoning on natural image-based Bongard Problems.
+We propose three structured paradigms—Direct Visual Rule Learning, Deductive Rule Learning, and Componential Analysis—designed to progressively enforce step-wise reasoning and disentangle the interplay between perception and reasoning.
+Our evaluation shows that advanced, closed-source VLMs (GPT-4o, Gemini 2.0) achieve near-superhuman performance, particularly when provided with high-quality image descriptions, while open-source models exhibit a significant performance bottleneck due to deficiencies in perception.
+An ablation study further confirms that perception, rather than reasoning, is the primary limiting factor, as open-source models apply extracted rules effectively when given accurate descriptions.
+These findings underscore the critical role of robust multimodal perception in enhancing generalizable visual reasoning and highlight the importance of structured, step-wise reasoning paradigms for advancing machine intelligence.
 Cognitive Paradigms for Evaluating VLMs on Visual Reasoning Task
 1 Introduction
-Human intelligence seamlessly integrates visual perception with language and higher-level cognition, enabling reasoning about visual information Kunda (2020).
-This ability, known as visual reasoning, supports tasks such as understanding spatial relationships, inferring object functions, and grounding language in the physical world.
+Human intelligence seamlessly integrates visual perception with language and higher-level cognition, enabling sophisticated reasoning about the visual world Kunda (2020).
+This ability, known as visual reasoning, supports tasks ranging from understanding spatial relationships to inferring object functions.
 Despite advancements in natural language processing (NLP) Zhao et al.
-(2023), achieving robust visual reasoning in machines remains a significant challenge, particularly in multimodal contexts Cao et al.
+(2023), achieving robust visual reasoning in machines, especially in complex, multimodal contexts, remains a significant challenge Cao et al.
 (2024).
-Large Language Models (LLMs) have transformed NLP by excelling in tasks like machine translation, text generation, and question answering Brown et al.
+Large Language Models (LLMs) have revolutionized NLP, demonstrating impressive capabilities in tasks like machine translation, text generation, and question answering Brown et al.
 (2020); Devlin (2018).
@@
 (2021) and logical reasoning Wei et al.
-(2022) suggest potential for bridging language and vision to tackle complex visual reasoning tasks.
-Visual reasoning tasks range from basic object recognition Lin et al.
-(2014) to abstract inference Zerroug et al.
+(2022a) suggest their potential for tackling complex visual reasoning tasks.
+However, effectively bridging language and vision to achieve human-like visual reasoning requires careful consideration of how these models process and interpret visual information.
+Visual reasoning challenges span a spectrum from basic object recognition Lin et al.
+(2014) to abstract inference Barrett et al.
+(2018); Zhang et al.
 (2019); Teney et al.
-(2020); Barrett et al.
-(2018); Zhang et al.
+(2020); Zerroug et al.
 (2022).
-Complex tasks demand integrating perceptual data with higher-order cognition, such as spatial reasoning and relational understanding Cao et al.
-(2024); Vaishnav et al.
+One such challenge – Bongard Problems (BPs) Bongard (1968), involving abstract rules distinguishing two image sets, exemplify the challenges of integrating perceptual data with higher-order cognition.
+They require both perceptual and conceptual reasoning, making them a valuable testbed for evaluating visual reasoning capabilities.
+Vision-Language Models (VLMs), such as Phi-1.5 Javaheripi et al.
+(2023), Llama-vision Meta (2024), Gemini 2.0 Google (2024), and GPT-4o OpenAI (2024), offer a promising avenue for addressing BPs.
+By leveraging multimodal training data, these models combine visual perception with higher-order cognitive processes.
+However, the mechanisms by which VLMs achieve visual reasoning, and how these mechanisms align with human cognitive strategies, remain an open question.
+This paper investigates the visual reasoning abilities of VLMs on natural image-based Bongard Problems using the Bongard Openworld Dataset Wu et al.
 (2024).
-Bongard Problems (BPs) Bongard (1968), involving abstract rules distinguishing two image sets, exemplify such challenges by requiring both perceptual and conceptual reasoning.
-Structured reasoning tasks like BPs require the reasoning process to follow specific rules or structures.
-These tasks involve identifying patterns or relationships among image components and applying abstract reasoning based on those patterns.
-Vision-Language Models (VLMs), such as Phi-1.5 Javaheripi et al.
-(2023), Llama-vision Meta (2024), Gemini Google (2024), and GPT-4o OpenAI (2024), provide a robust foundation for tackling BPs.
-By leveraging multimodal datasets, these models combine visual perception with higher-order cognitive processes, making them well-suited for structured reasoning tasks like BPs.
-This capability allows them to approach complex visual reasoning by breaking down information into identifiable, logical steps.
-Figure 1: Illustration of a Bongard OpenWorld dataset showing three positive and negative samples along with a query image.
-Distinguishing rule in this example is A group photo at a wedding reception..
-Based on the rule, query image belongs to negative category.
-This paper investigates the ability of Vision-Language Models (VLMs) to solve natural image-based Bongard Problems using the Bongard Openworld Dataset Wu et al.
-(2024) (example shown in Figure 1).
-Inspired by human problem-solving strategies, we evaluate three paradigms: holistic analysis, deductive rule learning, and componential analysis.
-Our findings demonstrate that VLMs employing these strategies can surpass human performance on this challenging visual reasoning task.
-Additionally, through ablation studies and a comparative analysis of open- and closed-source models, we identify key limitations in multi-image reasoning, such as difficulties with fine-grained distinctions, synthetic images, and nuanced contextual understanding.
-These insights provide critical directions for advancing the robustness and human-like reasoning capabilities of VLMs.
+Inspired by human problem-solving strategies, we introduce three distinct paradigms—direct visual rule learning, deductive rule learning, and componential analysis—designed to dissect the interplay between perception and reasoning in VLMs.
+Our key contribution is the development and application of these cognitively-inspired paradigms, which allow us to isolate and analyze the distinct roles of perception and reasoning in VLM performance.
+We demonstrate that advanced VLMs can achieve near- and even super-human performance on this challenging task, but also reveal critical limitations in the perception modules of open-source VLMs.
+These findings offer valuable insights into the computational underpinnings of visual reasoning and provide crucial directions for future research.
 Research Contributions:
   1. 1.
-    We evaluate the effectiveness of Vision-Language Models (VLMs) on the challenging task of solving Bongard Openworld Dataset—natural image-based BPs.
+    Three distinct, cognitively-inspired paradigms for evaluating VLM’s visual reasoning ability: Direct Visual Rule Learning, Deductive Rule Learning, and Componential Analysis, isolating the roles of perception and reasoning.
   2. 2.
-    We propose and evaluate three distinct paradigms for applying VLMs to this task, inspired by human learning and problem-solving strategies:
-    1. 2.1.
-      Holistic Analysis (Analyzing all information together)
-    2. $2.$
-      Deductive Rule Learning (Focusing on deriving rules before testing scenarios)
-    3. 2.3.
-      Componential Analysis (Decomposing images into structured, interpretable components)
+    Demonstration of near- and super-human performance by VLMs, particularly with high-quality image descriptions.
   3. 3.
-    We demonstrate that VLMs can surpass human performance in this challenging open-world visual reasoning task using human-centric learning methodologies.
+    Identification of a critical bottleneck in open-source VLMs: the perception module (image description generation), which significantly limits their reasoning capabilities.
   4. 4.
-    We provide a detailed evaluation of both open-source and closed-source models, highlighting current limitations and performance discrepancies in structured reasoning tasks.
-  5. 5.
-    We demonstrate that addressing limitations in fine-grained distinctions, synthetic image processing, and contextual understanding is key to improving VLM robustness and aligning them more closely with human reasoning capabilities.
+    Highlighting the crucial role of robust image descriptions for complex visual reasoning and the need for future research to focus on improving the perception components of open-source VLMs.
 2 Related Work
-To evaluate the capabilities of VLMs, various benchmarks have been developed, focusing on different aspects of visual understanding and reasoning.
-These benchmarks play a crucial role in assessing model performance and identifying areas for improvement.
-Traditional benchmarks focussed on single image tasks like image captioning Lin et al.
+Evaluating Vision-Language Models (VLMs) requires diverse benchmarks targeting various aspects of visual understanding and reasoning.
+Traditional benchmarks focused on single-image tasks like image captioning Lin et al.
 (2014), OCR Mishra et al.
@@
 (2018).
-For multi-image comprehension, researchers have developed interleaved image-text corpora, including MMC4 Zhu et al.
+However, the field has increasingly shifted towards more complex scenarios, including multi-image comprehension.
+This shift has led to the development of interleaved image-text corpora like MMC4 Zhu et al.
 (2024) and OBELICS Laurençon et al.
 (2024); Lin et al.
-(2024), along with dedicated benchmarks like Q-Bench Wu et al.
+(2024), along with dedicated benchmarks specifically designed for multi-image understanding, such as Q-Bench Wu et al.
 (2023), Memontos Wang et al.
@@
 (2023).
-While these benchmarks have contributed to progress in evaluating multi-image understanding, they often lack comprehensive coverage of complex reasoning scenarios.
-Other benchmarks like MMIU (Multimodal Multi-Image Understanding) Meng et al.
-(2024) have been pivotal in evaluating VLMs comprehensively, incorporating diverse datasets covering areas like science and mathematics.
-Despite advancements in VLM benchmarks that involve language, purely visual reasoning tasks without linguistic input remain crucial for evaluating core visual understanding.
-Bongard problems Bongard (1968) present a unique challenge in this regard.
-Their significance lies in their focus on abstract visual concepts, the absence of explicit language, and the inherent multi-image relational reasoning required for their solution.
-Extensions such as Bongard-HOI Jiang et al.
+Benchmarks like MMIU Meng et al.
+(2024) further broaden the scope by incorporating diverse datasets.
+While these benchmarks represent significant progress, they often fall short in capturing the intricacies of abstract visual reasoning, particularly those devoid of linguistic input.
+Purely visual reasoning tasks, where language is absent, are crucial for isolating and evaluating core visual understanding.
+Bongard problems Bongard (1968) provide an excellent example of such tasks, focusing on abstract visual concepts and requiring multi-image relational reasoning.
+Datasets like Bongard-HOI Jiang et al.
 (2022) and Bongard-OpenWorld Wu et al.
-(2024), further extend these challenges to human-object interactions and real-world scenarios, highlighting models’ few-shot reasoning capabilities.
-Beyond Bongard problems, other purely visual reasoning challenges have been proposed, each focusing on different aspects of visual cognition.
-These include Synthetic Visual Reasoning Test (SVRT) Fleuret et al.
-(2011), Compositional Visual Reasoning (CVR) Zerroug et al.
+(2024) extend these challenges to human-object interactions and real-world scenarios, respectively, pushing models to demonstrate robust few-shot reasoning capabilities.
+This focus on purely visual reasoning is not unique to Bongard problems.
+Other notable benchmarks in this category include SVRT Fleuret et al.
+(2011), CVR Zerroug et al.
 (2022), Raven’s Progressive Matrices (RPM) Barrett et al.
 (2018); Zhang et al.
-(2019), Visual Progressive Matrices (V-PROM) Teney et al.
-(2020), and the Abstract Reasoning Corpus (ARC) Chollet (2019).
-These benchmarks collectively assess a range of visual reasoning abilities, including spatial reasoning, compositional reasoning, and abstract pattern recognition, providing valuable insights into the visual understanding capabilities of AI models.
-Human cognitive processes offer valuable insights for evaluating visual reasoning in VLMs.
-Three primary learning strategies are particularly relevant: holistic analysis, deductive analysis, and componential analysis.
-Holistic analysis involves integrating all available information simultaneously, akin to similarity-based strategies where new information is compared to existing knowledge von Helversen et al.
-(2014); Johnson-Laird (2010).
-This approach emphasizes understanding the overall context and interconnections between elements.
-Focused problem-solving prioritizes understanding underlying principles through deductive reasoning before application, mirroring rule-based strategies Knauff and Wolf (2010); Helie et al.
-(2015); Rips (1994).
-This strategy emphasizes abstract rule extraction and application.
-Finally, componential analysis breaks down problems into individual components, extracting relevant details before integration Wu and Molnár (2022); Gluck et al.
+(2019), V-PROM Teney et al.
+(2020), and ARC Chollet (2019).
+These benchmarks collectively assess a wide range of visual reasoning abilities, including spatial, compositional, and abstract reasoning, offering valuable insights into the visual cognition of AI models.
+Our work is also informed by research in human cognition, which has identified key learning strategies relevant to visual reasoning.
+We consider three key learning strategies that are relevant to visual reasoning: direct visual rule learning – integrating all information simultaneously, akin to similarity-based strategies von Helversen et al.
+(2014); Johnson-Laird (2010), deductive reasoning – prioritizing rule extraction and application Knauff and Wolf (2010); Helie et al.
+(2015); Rips (1994), and componential analysis – breaking down problems into individual components Wu and Molnár (2022); Gluck et al.
 (2008); Hutchinson (1992).
-This approach emphasizes detailed analysis of individual elements and their subsequent combination.
-These distinct human learning strategies provide a framework for designing and interpreting visual reasoning benchmarks for VLMs, allowing researchers to assess how well these models emulate different aspects of human cognition.
+These strategies inform the design and interpretation of our visual reasoning benchmarks, enabling us to assess how well VLMs emulate different facets of human cognition.
+Finally, our work has connections to the growing body of research on Chain-of-Thought (CoT) prompting Kojima et al.
+(2022); Wei et al.
+(2022b); Huang and Chang (2023).
+While our focus is on visual reasoning and human cognitive strategies, CoT prompting, which encourages models to generate intermediate reasoning steps, shares a similar goal of improving model reasoning abilities.
+Recent work has explored applying CoT to multimodal tasks Zhang et al.
+(2023); Chen et al.
+(2024); Shao et al.
+(2024), demonstrating its potential for enhancing performance.
+Our work, by explicitly separating to visual modality alone focusing on perception and reasoning and analyzing the models’ ability to extract and apply rules, provides a complementary perspective on understanding and improving visual reasoning in VLMs.
 3 Models
@@
 This diversity allowed us to assess the strengths and limitations of different model types and scales.
-This selection encompasses state-of-the-art multimodal models, including GPT-4o OpenAI (2024), Gemini-2.0 Google (2024), Pixtral-12B Agrawal et al.
+Our selection encompasses state-of-the-art multimodal models, including GPT-4o OpenAI (2024), Gemini 2.0-flash-exp, henceforth Gemini 2.0 Google (2024), Pixtral-12B Agrawal et al.
 (2024), Llama-Vision-3.2 (11B, and 90B) Meta (2024), LLaVA (7B, 13B and 34B) Liu et al.
 (2023), and LLaVA built on Llama 3 (8B) XTuner (2025).
-By leveraging this variety, we systematically analyzed model performance across different architectural designs and parameter scales, offering a holistic view of their capabilities.
-For the evaluation purpose, two metrics are used: classification accuracy and semantic similarity (more in Appendix C).
-Further technical details, including model architectures and individual summaries, are elaborated in Appendix D.
+By leveraging this variety, we systematically analyzed model performance across different architectural designs and parameter scales, offering a holistic view.
+For the evaluation purpose, two metrics are used: classification accuracy and semantic similarity.
+Further technical details, including model architectures, evaluation and individual summaries, are elaborated in Appendix A.4.
 4 Dataset
@@
 (2024), a benchmark designed to assess few-shot visual reasoning abilities.
-The full dataset contains 1001 cases, from which we have chosen a sample of 500 cases for our experiments.
-Each case in the dataset consists of six positive and six negative examples and a query image.
-These cases are governed by an underlying "commonsense" rule, with the rules themselves categorized into distinct visual concepts (further details can be found in Appendix B).
-Figure 1 illustrates the task: given a set of positive and negative examples, the goal is to determine the underlying rule and classify a query image accordingly.
+The full dataset contains 1001 cases, from which we have chosen a sample of equally balanced 500 test samples for our experiments (further details can be found in Appendix A.3).
+Each case consists of six positive and six negative examples and a query image.
+These cases are governed by an underlying “commonsense” rule, with the rules themselves categorized into distinct visual concepts.
+Figure A.1 illustrates an example task where the identifying rule is to detect a group wedding photo.
 5 Experiments
 This paper investigates VLMs’ ability to solve natural image-based Bongard Problems.
-We evaluate three paradigms: holistic analysis (processing all information together), deductive rule learning (deriving rules first), and componential analysis (examining components separately).
-5.1 Holistic Analysis
-In this section, we describe our experiments using a holistic analysis approach, where the model is presented with all images (positive, negative, and test) simultaneously.
-The holistic analysis experiments assess the ability of VLMs to directly infer patterns and rules from a set of images.
-We designed a prompt that presents the model with m positive examples, n negative examples, and a final query image.
-The prompt instructs the model to:
-  1. 1.
-    Identify a rule that distinguishes the positive examples from the negative examples;
-  2. 2.
-    Analyze the query image;
-  3. 3.
-    Classify the query image as either positive or negative based on the identified rule.
-It is explicitly requested to the VLM model to provide its analysis, the derived rule, a description of the query image, and the final classification category. Details are provided in Appendix E.1.
+We evaluate three paradigms: direct visual rule learning (processing all information together), deductive rule learning (deriving rules first), and componential analysis (examining components separately).
+For each paradigm, the VLMs were prompted to provide four distinct outputs: (1) an overall analysis of the visual features and their relationships, (2) the derived rule that distinguishes the positive and negative categories, (3) a description of the query image, and (4) the final classification category for the query image.
+This consistent, structured output format facilitated a detailed examination of the models’ reasoning processes across all paradigms.
+Figure 1: A Multi-Stage Evaluation Framework for Visual Reasoning.
+Direct Visual Rule Learning: Positive and negative examples, along with a query image, are processed concurrently to identify the distinguishing rule and categorize.
+Deductive Rule Learning: A rule is extracted from the positive and negative examples and then applied to classify the query image.
+Componential Analysis: Textual descriptions of the images are generated, and rule extraction and classification are performed based solely on these textual representations.
+5.1 Direct Visual Rule Learning
+We begin by exploring a holistic approach to visual reasoning.
+In this paradigm, the Vision-Language Model (VLM) is presented with all images relevant to a given Bongard problem – the positive examples, the negative examples, and the query image – simultaneously.
+This approach aims to mimic human problem-solving strategies where all available information is considered concurrently.
+The prompt instructs the model to: (1) provide the overall analysis, (2) identify a rule that distinguishes the positive examples from the negative examples; (3) analyze the query image; (4) classify the query image as either positive or negative based on the identified rule The prompt is detailed in Appendix A.5.1.
 5.2 Deductive Rule Learning
-This section describes our experiments using a deductive rule learning approach, which involves a two-step process: first, the model infers a general rule; second, it applies this rule to classify a query image.
-The deductive rule learning experiments assess the ability of VLMs to explicitly formulate a rule before classifying a query image.
-This approach mirrors human deductive reasoning.
-We employed a two-stage prompting strategy.
-In the first stage, we presented the model with m positive examples and n negative examples and asked it to identify a rule that distinguishes the two categories.
-The prompt requested a concise summary of the identified rule (less than 20 words).
-The precise wording of this first-stage prompt is provided in Appendix E.2.
-In the second stage, we presented the model with the previously generated rule summary and a query image.
-The prompt instructed the model to analyze the query image in the context of the given rule and classify it as either positive or negative.
-The prompt also requested a brief analysis of the query image and a restatement of the rule.
-The precise wording of this second-stage prompt is provided in Appendix E.2.
+A core aspect of human intelligence is the ability to deduce general rules from specific examples and then apply those rules to new situations.
+This paradigm emulates human deductive reasoning, involving two key steps: first, inferring a general rule from a set of positive and negative examples; second, applying this derived rule to classify a query image.
+To investigate this capability in VLMs, we designed a two-stage prompting strategy.
+In the first stage, the model was presented with six positive examples and negative examples.
+The prompt, detailed in Appendix A.5.2.1, instructed the model to identify a distinguishing rule that separates the two categories.
+Critically, the model was asked to provide a concise summary of this identified rule, limited to a maximum of 20 words.
+This constraint encouraged the model to focus on the most salient features distinguishing the categories.
+In the second stage, the model was provided with the previously generated rule summary and a query image.
+The prompt, detailed in Appendix A.5.2.2, instructed the model to analyze the query image in the context of the given rule and classify it as either positive or negative.
+The model was also asked to provide a brief analysis of the query image and restate the rule it was applying.
+This allowed us to examine not only the classification decision but also the model’s reasoning process.
 5.3 Componential Analysis
-This section describes our experiments using a componential analysis approach (detailed in Appendix E.3).
-This method involves decomposing each image into its constituent visual elements and then reasoning based on these components (textual elements).
-This process is divided into three main stages: (1) detailed image description generation, (2) providing instruction on rule derivation, and (3) rule extraction and application based on these descriptions.
-The componential analysis experiments aim to evaluate the ability of VLMs to reason based on structured descriptions of visual content.
-This approach simulates how humans might break down complex visual scenes into simpler parts for analysis.
-We employed a three-stage prompting strategy.
-In the first stage, we generate detailed descriptions of each image.
-This prompt instructs the model to analyze an image and provide a structured description, covering aspects such as scene, objects (living and inanimate), activities, contextual elements, visual patterns, emotional undertones, textual information, and a summary.
-In the second stage, we provide the models with instructions on how to analyze the outputs and derive the rule.
-In the final stage, which is purely textual, we provided the model with the descriptions generated in the first stage, along with information about which images belong to positive examples and negative examples, and the query image.
-The prompt then asks the model to derive a rule based on the descriptions of positive and apply this rule to classify the query image.
-5.4 Rule-Based Evaluation
-This section describes experiments designed to disentangle the image analysis (perception) and rule application (abstract reasoning) capabilities of VLMs.
-Unlike the previous experiments, where the models were tasked with both extracting and applying rules, here we provide the models with pre-extracted rules and assess their ability to apply these rules to classify query images.
-These experiments isolate the rule application component of the Bongard problem-solving process.
-This experimental setup is conceptually similar to the Deductive Rule Learning experiments (Section 5.2), which also involved a two-step process: (1) inferring a general rule and (2) applying that rule to classify a query image.
-However, in this experiment, we bypass the first step (rule inference) and instead use the rule summaries generated by GPT-4o in the Deductive Rule Learning experiment.
-For each Bongard problem, we took the rule summary generated by GPT-4o.
-We then presented the model with the query image and the corresponding rule summary, explicitly instructing it to classify the query image based on the provided rule.
-This approach allows us to specifically assess how well the models can understand and apply a given rule, independent of their ability to extract the rule themselves.
+While deductive rule learning operates at the level of abstract rules, componential analysis takes a more granular approach.
+This method decomposes each image into its constituent visual elements and reasons based on these components, represented textually.
+This approach mirrors how humans often break down complex visual scenes into simpler parts for analysis.
+The process unfolds in two distinct stages: (1) detailed image description generation and (2) rule derivation and application.
+The componential analysis experiments aim to evaluate the ability of VLMs to reason based on structured, textual descriptions of visual content.
+We employed a two-stage prompting strategy to simulate this compositional reasoning process.
+In the first stage, detailed descriptions of each image (positive, negative, and query) were generated using the prompt detailed in Appendix A.5.3.1.
+This prompt instructs the model to analyze an image and provide a structured textual description covering aspects such as scene, objects (living and inanimate), activities, contextual elements, visual patterns, emotional undertones, textual information, and a summary.
+In the second stage, the model was provided with the descriptions generated in the first stage, along with explicit “textual” information about which images correspond to positive examples and which correspond to negative examples, as well as the description of the query image.
+The prompt, detailed in Appendix A.5.3.2, then instructed the model to derive a rule based on the descriptions and apply this rule to classify the query image – described in textual form.
+This stage combines the processes of rule derivation and application, relying solely on the textual representations of the visual content.
+This isolates the models’ ability to reason about visual concepts based on their textual representations.
 6 Results
-6.1 Holistic Analysis
-Holistic analysis necessitates models with multi-image and multimodal capabilities. Among the models tested, only Gemini 2.0 (flash) (referred to Gemini here) and GPT-4o possess these features.
+6.1 Direct Visual Rule Learning
+Direct visual rule learning necessitates models with multi-image and multimodal capabilities. Among the models tested, only Gemini 2.0 and GPT-4o possess these features.
 Table 1 presents the overall accuracy, as well as the accuracy on positive and negative samples separately.
-As shown, Gemini achieves a slightly higher overall accuracy of 82.2% compared to GPT-4o’s 80.0%.
-Unlike GPT-4o, Gemini shows a more balanced performance across both positive and negative samples, suggesting a more robust approach to handling cases where no clear rule might exist for the negative examples.
+As shown, Gemini 2.0 achieves a slightly higher overall accuracy of 82.2% compared to GPT-4o’s 80.0%.
 These results suggest that while GPT-4o may excel at recognizing patterns within the positive examples, its lower accuracy on negative samples indicates challenges in distinguishing subtle counter examples.
-On the other hand, Gemini’s balanced performance reflects greater robustness in addressing the diverse reasoning demands of holistic analysis tasks.
-Model Holistic Analysis Deductive Rule Learning Componential Analysis
+On the other hand, Gemini 2.0’s balanced performance reflects greater robustness in addressing the diverse reasoning demands of direct visual rule learning tasks.
+6.2 Deductive Rule Learning
+The deductive rule learning paradigm, with its two-stage process of rule extraction and application, offers a more structured approach to visual reasoning compared to the direct visual rule learning.
+This structured approach allows us to evaluate the model’s ability to not only identify patterns but also articulate them as explicit rules, providing insights into its deductive reasoning capabilities.
+We conducted these experiments using Gemini 2.0 and GPT-4o.
+Table 1 summarizes the overall accuracy, as well as the accuracy on positive and negative samples separately.
+As shown in Table 1, GPT-4o achieves slightly higher overall accuracy than Gemini 2.0, suggesting greater effectiveness in combining rule extraction and application within this paradigm.
+While both models demonstrate relatively balanced performance across positive and negative samples, subtle differences emerge.
+Similar to the previous experiment, GPT-4o exhibits a stronger ability to classify positive samples, and Gemini 2.0 achieves more consistent performance across both categories, suggesting greater robustness across diverse reasoning scenarios.
+Comparing these results to those obtained using the direct visual rule learning (Table 1), we can see a clear improvement in overall accuracy for both models.
+This indicates that the two-stage deductive approach enhances the models’ ability to handle the complexities of both positive and negative test samples, surpassing the performance seen in the holistic approach.
+This improvement likely stems from the explicit separation of the rule extraction and application processes, allowing the models to focus on each task individually.
+This suggests that explicitly structuring the reasoning process, even within purely visual tasks, benefits VLMs.
+Model Direct Visual Rule Learning Deductive Rule Learning Componential Analysis
 neg pos Overall neg pos Overall neg pos Overall
 GPT-4o 66.4 93.6 80.0 82.8 93.2 88.0 92.8 92.8 92.8
-Gemini 2.0 80.48 83.94 82.2 85.66 87.95 86.8 90.84 96.39 93.6
+Gemini 2.0 80.5 83.9 82.2 85.7 87.9 86.8 90.8 96.4 93.6
 Pixtral-12B - - - - - - 59.2 48.4 53.8
-Llava-Llama3-8B - - - - - - 93.02 13.08 53.15
+Llava-Llama3-8B - - - - - - 93.0 13.1 53.2
 Llama-vision-11B - - - - - - 42.8 64.0 53.4
-Llava 7B - - - - - - 65.20 67.20 66.20
+Llava 7B - - - - - - 65.2 67.2 66.2
 Human - - - - - - - - 91.0
 Table 1: Classification accuracy across different Analysis Paradigms and Human performance Wu et al. (2024). neg and pos represents the classification accuracy for negative and positive query images.
-6.2 Deductive Rule Learning
-This two-stage approach allows us to evaluate the model’s ability to not only identify patterns but also articulate them as explicit rules, providing insights into its deductive reasoning capabilities.
-We conducted these experiments using Gemini and GPT-4o.
-Table 1 summarizes the overall accuracy, as well as the accuracy on positive and negative samples separately.
-This time GPT-4o achieves a slightly higher overall accuracy compared to Gemini.
-Both models demonstrate balanced performance across positive and negative samples.
-GPT-4o, in particular, excels in classifying positive samples, while Gemini achieves a more consistent performance across both categories.
-These results indicate that the two-stage deductive approach enhances the models’ ability to handle the complexities of both positive and negative examples, surpassing the performance seen in the holistic approach.
-GPT-4o’s superior accuracy on positive samples suggests its effectiveness in identifying and applying explicit patterns, while Gemini’s balanced results highlight its robustness across diverse reasoning scenarios.
 6.3 Componential Analysis
-The componential analysis approach provides a framework for evaluating a broader spectrum of vision-language models (VLMs) by decomposing the complex multi-image reasoning task into manageable single-image description tasks.
-This design facilitates the inclusion of models that lack explicit multi-image processing capabilities but can effectively analyze individual images and process structured textual representations.
-We conducted the three-stage componential analysis experiment using six VLMs: Pixtral-12B, Llava-Llama3-8B, Llama-Vision-11B, Llava 7B, GPT-4o, and Gemini.
-The performance metrics, presented in Table 1, include overall accuracy and accuracy on positive and negative samples.
-The results demonstrate a wide range of performance across the evaluated models.
-GPT-4o and Gemini achieve near-optimal accuracies, with overall accuracies of 92.8% and 93.6%, respectively.
-Notably, these results surpass human-level performance, which is reported at 91% on this task.
-This finding highlights the potential of advanced VLMs to excel in tasks traditionally dominated by human reasoning capabilities.
-Open-source models, including Pixtral-12B, Llava-Llama3-8B, and Llama-Vision-11B, Llava 7B exhibit significant variability.
-Llava-Llama3-8B, for example, achieves a high accuracy on negative samples but an exceptionally low accuracy on positive samples, indicating a strong bias toward negative classifications.
-Similarly, Llama-Vision-11B demonstrates an inverse pattern.
-These discrepancies suggest that these models struggle to generalize effectively across different classes.
-The superior performance of GPT-4o and Gemini indicates that structured, detailed descriptions significantly enhance reasoning capabilities in VLMs.
-Their ability to surpass human-level performance underscores the importance of integrating multimodal capabilities with robust processing pipelines.
-Furthermore, performance open-source models highlights the potential for continued advancements in accessible VLMs to achieve substantial progress in complex reasoning tasks.
+The componential analysis approach evaluates VLMs by decomposing the complex multi-image reasoning task into manageable single-image description tasks.
+This method separates the perception component (image analysis and description) from the reasoning component (rule derivation and application), relying solely on textual representations for the latter.
+This isolation of perception and reasoning allows for a deeper understanding of their respective contributions to VLM performance.
+Table 1 presents the performance metrics for six VLMs: Pixtral-12B, Llava-Llama3-8B, Llama-Vision-11B, Llava 7B, GPT-4o, and Gemini 2.0.
+The results reveal significant performance variations, highlighting the crucial role of accurate image descriptions in complex visual reasoning.
+GPT-4o and Gemini 2.0 achieve near-optimal accuracies (92.8% and 93.6%, respectively), surpassing human-level performance i.e.
+91% Wu et al.
+(2024).
+This demonstrates the potential of advanced VLMs when provided with high-quality, structured descriptions.
+Their strong performance suggests that these models excel at leveraging textual information for rule derivation and application—the reasoning component.
+In contrast, the open-source models exhibit greater performance variability.
+While Pixtral-12B and Llava 7B achieve moderate, balanced performance, Llava-Llama3-8B and Llama-Vision-11B show a striking bias: one performs well on negative samples but poorly on positive samples, while the other exhibits the opposite trend.
+This suggests a significant challenge in generalization for these models, likely stemming from inconsistencies or inaccuracies in their generated descriptions—the perception component.
+The stark difference in performance between the closed- and the open-source models underscores the importance of robust image description generation.
+The near- or super-human performance of GPT-4o and Gemini 2.0 highlights the potential of VLMs when both perception and reasoning components are strong.
+The challenges faced by the open-source models, suggest that future research should prioritize improving the perception modules of these models to achieve human-like reasoning.
+Further investigation is needed to understand the specific factors contributing to the open-source models’ difficulties in generating consistent and accurate descriptions.
+7 Ablation
 7.1 Rule-Based Evaluation
-The Rule-Based Evaluation experiment isolates the models’ ability to apply explicitly provided rules to classify query images, bypassing the rule inference step.
-This approach offers a focused assessment of the rule application capabilities of the models.
-The results, presented in Table 2, reveal distinct trends across the evaluated models.
-Pixtral-12B achieved the highest overall accuracy (83.8%), indicating a strong ability to comprehend and operationalize the provided rules effectively.
-Its relatively balanced performance on both negative and positive samples highlights its robust application of rules across varying contexts.
+The Deductive Rule Learning paradigm (Section 6.2) combines rule extraction and application.
+To isolate the latter, we designed an ablation study focusing solely on rule application, thus disentangling perception from reasoning.
+This allows us to directly assess the models’ reasoning capabilities independent of their perception modules.
+Specifically, we used the rule summaries generated by GPT-4o during the rule extraction phase.
+For each Bongard problem, we provided the open-source models with two inputs: 1) the corresponding rule summary, and 2) the query image.
+The models were instructed to classify the query image based on the provided rule extracted using GPT-4o.
+This isolates the models’ ability to understand and apply a given rule, independent of their capacity to extract it.
+Table 2 presents the results.
+Pixtral-12B achieved the highest overall accuracy (83.8%), demonstrating strong rule comprehension and application.
+Its balanced performance suggests robust rule application across contexts, contrasting with some other models exhibiting a bias towards confirming positive matches.
+While some achieve high positive accuracy, they struggle with negative cases, suggesting difficulty identifying when a rule does not apply.
+Pixtral-12B’s balanced performance suggests greater adeptness, while models like Llama-Vision-xx and LLaVA-xx may over-rely on confirmatory evidence.
+Comparing the performance of Pixtral-12B and the Llava models with their componential analysis results reveals a crucial insight.
+Pixtral-12B maintains relatively consistent performance across both evaluations.
+However, the Llava models’ accuracy drops substantially in the componential analysis compared to the rule-based evaluation.
+This performance gap strongly indicates that the Llava models’ primary bottleneck lies in the perception stage – the ability to generate accurate and comprehensive image descriptions – rather than in their reasoning capabilities (rule application), which are more effectively assessed when pre-extracted rules are provided.
+By isolating rule application, this ablation provides insights into the open-source models’ strengths and limitations in reasoning.
+Combined with the Deductive Rule Learning results, this analysis disentangles factors contributing to overall performance, distinguishing between deficits in rule extraction (a proxy for perception) and application (pure reasoning).
 Model Accuracy (%)
 Overall neg pos
-LLaVA-7B 66.20 65.20 67.20
+LLaVA-7B 66.2 65.2 67.2
 Llama-vision-11B 68.2 38.0 98.4
@@
 Table 2: Rule-Based Evaluation Results showing overall classification accuracy along with categorical accuracy for positive (pos) and negative (neg) samples.
-The results in Table 2 highlight that while certain models excel in confirming positive matches, their difficulties in handling negative cases expose potential biases in rule application mechanisms.
-Pixtral-12B’s relatively balanced performance suggests it is more adept at applying rules across diverse scenarios, while models like Llama-Vision-xx and LLaVA-xx may overfit to confirmatory evidence.
-By isolating the rule application component, this experiment provides valuable insights into the models’ strengths and limitations.
-When combined with findings from the Deductive Rule Learning experiments (Section 5.2), this analysis helps to disentangle the factors contributing to model performance, distinguishing between deficits in rule extraction and rule application.
-Such insights are crucial for guiding improvements in visual-language model reasoning capabilities.
-7 Ablation
+7.2 Analysis per Commonsense Category
+To better understand the comparative capabilities of GPT-4o and Gemini 2.0, we analyzed their performance across ten commonsense reasoning categories – provided in the Bongard Openworld dataset (details in Appendix A.3.3).
+This analysis highlights the models’ strengths and limitations in handling diverse commonsense reasoning categories (refer to Table A.3).
+Both models demonstrated robust performance across most categories indicating models’ capability to process high-level conceptual patterns effectively.
+Differences in performance were observed in categories requiring more nuanced or contextual understanding, such as “Taste / Nutrition / Food” and “Human-Object Interaction.” This suggests that while both models are adept at general reasoning tasks, their architecture and training data influence their performance on specific types of visual information.
+Notably, Gemini 2.0 showed strengths in scenarios involving fine-grained object attributes, while GPT-4o performed better in tasks demanding contextual interpretation.
 7.3 Semantic Similarity Analysis
-To further analyze the reasoning process of the best-performing models, GPT-4o and Gemini, we conducted a semantic similarity analysis.This analysis aims to quantify the alignment between the descriptions of the query images and the rules identified for each category.
-We hypothesized that for correctly classified positive examples, the similarity between the description and rule should be high, while for negative examples, it should be lower.
-[…]
-7.2 Analysis per Commonsense Category
-To better understand the comparative capabilities of GPT-4o and Gemini, we analyzed their performance across ten commonsense reasoning categories in the Bongard Openworld dataset (details in Appendix B.3).
-This analysis (in Table A.5) highlights the models’ strengths and limitations in handling diverse commonsense reasoning categories.
-Both models demonstrated robust performance across most categories indicating models’ capability to process high-level conceptual patterns effectively.
-Differences in performance were observed in categories requiring more nuanced or contextual understanding, such as "Taste / Nutrition / Food" and "Human-Object Interaction." This suggests that while both models are adept at general reasoning tasks, their architecture and training data influence their performance on specific types of visual information.
-Notably, Gemini showed strengths in scenarios involving fine-grained object attributes, while GPT-4o performed better in tasks demanding contextual interpretation.
+To further analyze the reasoning process of the best-performing models, GPT-4o and Gemini 2.0, we conducted a semantic similarity analysis.
+This analysis aims to quantify the alignment between the perception (image description) and the reasoning component (identified rule).
+We hypothesized that for correctly classified positive examples, the similarity between the description and rule should be high, reflecting a strong connection between perception and the derived rule.
+Conversely, for negative examples, the similarity should be lower, as the description should not align with the rule.
+Using cosine similarity of sentence embeddings generated using OpenAI embeddings, we computed the mean and standard deviation of the semantic similarity scores for both models and categories.
+Both models exhibit higher similarity (see Table A.4) for positive examples compared to negative ones, confirming that correct classifications align with the identified rule.
+However, the small difference in similarity ($\sim$ 6%) suggests that the negative samples were designed to be closer to positive samples to make them challenging.
+This highlights the subtle perceptual distinctions required for correct classification in these challenging cases.
 7.4 Analysis of Incorrectly Classified Samples
-In this ablation study, we analyzed the misclassified samples from GPT-4o and Gemini during the Componential Analysis experiments to identify common limitations.
+Here we analyzed the misclassified samples from GPT-4o and Gemini 2.0 during the Componential Analysis experiments to identify common limitations.
 The total 15 common test samples were incorrectly identified by both the VLMs.
-The errors revealed several recurring issues (more details in Table A.7), including challenges with synthetic images that lack natural textures, difficulties in making fine-grained distinctions, and struggles with recognizing specific image types, such as satellite views.
-Additionally, the models showed limitations in understanding nuanced contextual information, as evidenced by misclassifications related to artistic styles and the broader scene context.
-These findings highlight that while the models excel in many areas, they still face challenges in handling complex visual details and contextual reasoning, suggesting areas for improvement in VLMs’ robustness and generalization.
+The errors revealed several recurring issues (more details in Table LABEL:tab:tasks).
+We found eleven cases of errors in rule extraction; more specifically, four cases of being too general, one case of a missed property, five cases of the missing required object, two cases of focusing on a wrong property.
+We also found three cases of weak similarity-based reasoning disregarding some parts of the rule.
+These findings highlight that while the models excel in many areas, they still face challenges in handling complex visual details and reasoning, suggesting areas for improvement in VLMs’ robustness and generalization.
 8 Discussion
-Our study investigates the application of Vision-Language Models (VLMs) to Bongard Open-World Problems (BPs), providing insights into the computational mechanisms underlying visual reasoning and their alignment with human cognitive strategies.
-We introduced three evaluation paradigms—Holistic Analysis, Deductive Rule Learning, and Componential Analysis—designed to mirror distinct human approaches to learning and problem-solving.
-These paradigms allow us to evaluate VLM performance and understand how these models relate to human cognition.
-A critical aspect of human scene understanding is the dynamic interaction between the observer and their environment Zelinsky (2013); Malcolm et al.
+This study investigates how VLMs tackle Bongard Problems (BPs), offering insights into their visual reasoning mechanisms and alignment with human cognitive strategies.
+We introduced three evaluation paradigms, each mirroring a distinct human problem-solving approach.
+These paradigms allow us to evaluate VLM performance and dissect the contributions of different cognitive processes to their visual reasoning abilities.
+Human scene understanding is an active, interactive process Zelinsky (2013); Malcolm et al.
 (2016); Gibson (2015).
-This interactionist view posits that understanding a scene is an active process shaped by behavioral goals, rather than a passive reception of sensory input.
-This active engagement with the visual world is directly relevant to how both humans and VLMs approach BPs.
-Rather than processing individual images in isolation, both humans and VLMs benefit from considering the relationships between images in a given set.
-Our Holistic Analysis paradigm specifically targets this relational aspect by requiring VLMs to process all images and descriptions simultaneously.
-This approach allows the model to capture the overall context and directly identify the distinguishing rule, akin to human holistic problem-solving, where the "gist" of a scene is rapidly perceived, guiding subsequent attention and actions Biederman (1987); Li et al.
+This interactionist perspective is key to understanding how both humans and VLMs approach BPs.
+Rather than passively receiving sensory input, both benefit from actively considering the relationships between images.
+Our Direct Visual Rule Learning paradigm directly probes this, requiring simultaneous processing of all images, mirroring holistic human problem-solving where a rapid “gist” perception guides further analysis Biederman (1987); Li et al.
 (2002).
-Such holistic processing provides a crucial initial understanding, which can then be refined through more detailed analysis.
-This process heavily relies on working memory to maintain the "gist" while further analysis unfolds Baddeley (2012).
-Following this initial holistic understanding, human visual information processing is further refined through attentional mechanisms that guide gaze and cognitive processing Posner (1980).
-While eye movements can reflect attentional shifts, our Deductive Rule Learning paradigm directs the VLMs’ attention on learning the underlying rule prior to evaluating query images.
-This mirrors human deductive reasoning, where a hypothesis is formed and then tested against new evidence.
-By focusing on rule inference before rule application, the VLM can store learned rules in long-term memory Squire (1992), allowing for efficient application to unseen instances.
-In addition to holistic and deductive strategies, humans often use analytical approaches to break down complex problems.
-The Componential Analysis paradigm reflects this tendency by decomposing a complex problem into smaller, more manageable components.
-In this paradigm, VLMs analyze positive and negative examples separately, extracting individual rules for each set before comparing them to identify the distinguishing factor.
-This process aligns with cognitive theories that distinguish between System 1 (fast, intuitive perception) and System 2 (slow, deliberate reasoning) Kahneman (2011).
-Object recognition and categorization, a System 1 process involving perceptual mechanisms Marr (2010), precedes the more deliberate System 2 reasoning required to compare and contrast extracted rules.
-This analytical decomposition facilitates a more precise understanding of the underlying rules.
-Our findings, which show that VLMs can surpass human performance on BPs when applying these human-inspired paradigms, underscore the importance of attention and memory mechanisms in visual reasoning.
-Attention mechanisms, critical for selecting relevant information, are implemented in VLMs through architectures such as those described in Bahdanau et al.
-(2016) and Vaswani et al.
-(2017).
-These mechanisms allow VLMs to selectively focus on relevant image features, similar to human gaze patterns and attentional shifts.
-Meanwhile, memory mechanisms are essential for encoding visual information, maintaining it during rule formulation, and retrieving it for comparison.
-This is evident in the VLMs’ ability to store learned rules (long-term memory) and maintain representations of images during analysis (working memory).
-The interplay between attention and memory is crucial for both human and artificial intelligence, providing deeper insights into the computational foundations of visual reasoning Vaishnav and Serre (2023).
-Moreover, our evaluation of both open-source and closed-source models highlights significant performance discrepancies, shedding light on the limitations of current VLMs in handling complex visual reasoning tasks.
-These limitations underscore the need for further model refinement, particularly in handling multi-image reasoning tasks and improving rule extraction capabilities.
-Finally, we demonstrate that effective reasoning in complex visual scenarios requires the complementary strengths of VLMs and Large Language Models (LLMs).
-While VLMs excel at extracting features from images, LLMs play a crucial role in higher-level reasoning tasks, indicating a promising direction for future research where the integration of both types of models can enable more sophisticated problem-solving in visual reasoning tasks.
-9 Limitation
-Despite the promising results of this study, several limitations must be considered.
-The Bongard Openworld dataset may not fully capture the complexity and diversity of real-world visual reasoning tasks, potentially limiting generalizability.
-Additionally, the reliance on pre-extracted rule summaries in some experiments restricts the evaluation of models’ complete reasoning capabilities, particularly in rule extraction.
-The scalability of large models, such as GPT-4o and Gemini, is also a concern due to their high computational demands, and their ability to generalize to noisy, real-world data remains uncertain.
-Furthermore, the opacity of model decision-making processes highlights the need for better interpretability, and the risk of overfitting to task-specific knowledge could hinder broader applicability.
-Lastly, our evaluation primarily focuses on accuracy and semantic similarity, which may not fully capture the nuances of model reasoning.
-These limitations highlight areas for future research and refinement in applying VLMs to complex reasoning tasks.
+This initial holistic understanding is then typically refined through focused attention Posner (1980).
+Our Deductive Rule Learning paradigm simulates this by prompting VLMs to first learn the underlying rule before evaluating query images (more in Appendix A.2).
+This mimics human deductive reasoning where a hypothesis is formed and then tested.
+Humans also frequently employ analytical strategies, breaking down complex problems.
+Our Componential Analysis paradigm reflects this, with VLMs analyzing positive and negative examples separately before comparison.
+This aligns with dual-process theories of cognition, where intuitive System 1 - perception, precedes deliberate System 2 - reasoning Kahneman (2011); Marr (2010).
+Critically, this paradigm allows us to isolate the perception component (image description) from the reasoning component (rule derivation and application), revealing the distinct roles these play in VLM performance.
+Importantly, our three paradigms can be seen as progressively more detailed instantiations of a Chain-of-Thought (CoT) approach (more details in Appendix A.6.3).
+Direct Visual Rule Learning implicitly requires holistic image relationship consideration, Deductive Rule Learning adds explicit rule extraction and Componential Analysis represents the most granular CoT, decomposing images into separate descriptions.
+This progressive increase in detail is mirrored by increased performance, suggesting that structuring visual reasoning benefits VLMs.
+Unlike typical CoT prompting, our approach uses structured visual and textual representations in multiple stages, offering a way to induce step-by-step reasoning.
+Our results show that VLMs can surpass human performance on BPs given high-quality, structured descriptions, notably in the Componential Analysis paradigm.
+This suggests that mitigating perceptual challenges allows VLMs to effectively use textual representations for sophisticated reasoning.
+However, the substantial performance gap between open- and closed-source models reveals critical limitations, especially in the perception component.
+The stark contrast between rule-based evaluation (bypassing perception) and Componential Analysis (requiring it) underscores this bottleneck.
+For instance, Llava models exhibited a significant performance drop when generating descriptions compared to using pre-extracted rules.
+9 Limitations
+While this work offers valuable insights, it’s crucial to acknowledge its limitations.
+Our evaluation relies on the Bongard OpenWorld dataset, which may not fully capture the diversity of real-world visual reasoning scenarios.
+Exploring more diverse datasets is essential for assessing generalizability.
+The Componential Analysis paradigm, while insightful, is inherently tied to the quality of generated image descriptions, which are often imperfect, especially for open-source VLMs.
+Furthermore, the computational demands of state-of-the-art models raise concerns about scalability and accessibility.
+Finally, our quantitative metrics may not fully capture the nuances of model reasoning.
+Future work should prioritize improving open-source VLMs’ perception modules, exploring more diverse datasets, and developing richer evaluation metrics to provide deeper insights into model reasoning.
+Integrating VLMs and LLMs to leverage their complementary strengths is a promising direction for achieving more human-like reasoning.
 10 Conclusion
-This work investigates the effectiveness of Vision-Language Models (VLMs) in solving the challenging Bongard Openworld dataset, a task requiring complex visual reasoning.
-We proposed three human-inspired paradigms—Holistic Analysis, Deductive Rule Learning, and Componential Analysis—and demonstrated that VLMs, particularly GPT-4o and Gemini, can surpass human-level performance.
-Our results highlight that VLMs excel when combining detailed image analysis with structured rule-based reasoning.
-The study also emphasizes the complementary roles of VLMs and Large Language Models (LLMs), with VLMs excelling at visual feature extraction and LLMs crucial for higher-level reasoning.
-These findings provide valuable insights into model design and evaluation, suggesting a promising direction for future research in complex visual reasoning tasks and real-world applications.
+This work presents a cognitively-inspired framework for evaluating the visual reasoning capabilities of Vision-Language Models (VLMs) on natural image-based Bongard Problems.
+By systematically dissecting the interplay between perception and reasoning, we introduce three structured paradigms—Direct Visual Rule Learning, Deductive Rule Learning, and Componential Analysis—that progressively enforce step-wise reasoning.
+Our findings reveal that advanced, closed-source VLMs (GPT-4o, Gemini 2.0) achieve near-superhuman performance, particularly when provided with high-quality image descriptions, whereas open-source models exhibit a significant performance gap due to deficiencies in their perception modules.
+The ablation study further highlights that perception, rather than reasoning, is the primary bottleneck, as open-source models can apply extracted rules effectively when given accurate descriptions.
+These results underscore the critical role of robust multimodal perception in advancing generalizable and interpretable visual reasoning models.
+Addressing this challenge requires enhancing visual representations, improving multimodal alignment, and designing architectures that enforce structured, step-wise reasoning akin to Chain-of-Thought prompting.
+Additionally, integrating cognitive principles from human problem-solving strategies could provide a pathway toward developing more robust, transparent, and human-aligned visual reasoning systems.
 Acknowledgments
@@
 Appendix
-Appendix A Broader Relevance
-Broader Relevance
-The insights gained from this work have broader implications beyond the specific task of solving the Bongard Openworld dataset.
-As the capabilities of Vision-Language Models (VLMs) continue to evolve, the methodologies and paradigms explored in this paper can contribute significantly to the development of models capable of more advanced reasoning and problem-solving in real-world applications.
-The ability to tackle complex visual reasoning tasks, such as the Bongard problems, is critical for a range of fields, from autonomous systems and robotics to medical imaging and beyond.
-One of the most significant implications of our work is the demonstration that VLMs, when integrated with structured reasoning paradigms, can outperform human-level performance in specific problem-solving scenarios.
-This opens up new avenues for the deployment of AI systems in environments that require a blend of visual perception and reasoning, such as dynamic real-world settings where an agent must continually adapt to new and diverse scenarios.
-By surpassing human-level performance in tasks that require complex visual analysis, VLMs could enable more robust and adaptive AI systems that can be deployed in areas ranging from automated decision-making to real-time problem solving.
-Furthermore, our work also highlights the critical gap between the reasoning abilities of VLMs vs Large Language Models (LLMs).
-A combination of both is essential is needed for complex reasoning tasks.
-While VLMs excel at understanding and processing visual inputs, LLMs bring the reasoning and linguistic abilities necessary to contextualize and apply rules based on these inputs.
-This integration is essential for advancing research in cognitive modeling, where understanding the interplay between perception, reasoning, and language is central.
-Our findings encourage further exploration of hybrid architectures that seamlessly blend visual and language models, potentially driving progress in areas such as visual question answering, image captioning, and multimodal learning.
-Finally, our detailed evaluation of both open-source and closed-source models contributes to the ongoing dialogue about the accessibility, fairness, and transparency of AI systems.
-By examining the performance discrepancies across various architectures, we not only illuminate the current limitations of VLMs but also provide actionable insights for future improvements.
-This work could inspire the development of more efficient and capable models, particularly in the open-source community, and contribute to efforts aimed at making advanced AI more accessible and usable for a wide range of applications.
-In summary, this work sets the stage for more advanced, human-like AI systems capable of solving complex visual reasoning tasks, with far-reaching applications in domains such as autonomous systems, healthcare, and cognitive modeling.
-The methodologies and findings outlined here provide a valuable foundation for future research and development in AI, pushing the boundaries of what is possible with multimodal learning and reasoning.
-Appendix B Dataset
-B.1 Bongard OpenWorld Dataset
-We utilize a subset of 500 cases from the Bongard OpenWorld dataset Wu et al.
+Appendix A.1 Broader Relevance
+This study offers insights with broader implications for the development of more robust and human-like AI systems.
+Our cognitive inspired evaluation paradigms — Direct Visual Rule Learning, Deductive Rule Learning, and Componential Analysis—provide valuable tools for assessing and understanding the strengths and limitations of Vision-Language Models (VLMs) on complex visual reasoning tasks.
+The insights gained from applying these paradigms to the Bongard Open-World Dataset extend beyond this specific task, contributing to the development of VLMs capable of advanced reasoning in real-world applications.
+Our key finding—that the perception module (image description) is a critical bottleneck for open-source VLMs—has significant implications for future research.
+Improving the ability of these models to generate accurate and comprehensive image descriptions is crucial for bridging the performance gap with closed-source models and unlocking their full potential for complex visual reasoning.
+The demonstration that advanced VLMs (GPT-4o and Gemini 2.0) can achieve near- or even super-human performance on BPs, particularly when provided with high-quality descriptions, underscores the potential of these models for tasks requiring sophisticated visual understanding.
+This reinforces the importance of continued research into multimodal architectures that effectively integrate robust perception with advanced reasoning capabilities.
+Finally, our comparative evaluation of open- and closed-source models contributes to the broader discussion about AI accessibility and transparency.
+By highlighting the performance disparities, we identify specific areas for improvement in open-source models, promoting the development of more accessible and capable AI systems.
+This work provides a foundation for future research aimed at bridging the gap between human and machine visual reasoning, paving the way for more robust and reliable AI in diverse real-world applications.
+Appendix A.2 Attention and Memory in Visual Reasoning
+While our study primarily focuses on the interplay between perception and reasoning, the roles of attention and memory are also implicitly present in our paradigms.
+The Direct Visual Rule Learning paradigm, by requiring simultaneous processing of all images, likely engages VLMs’ “visual attention” mechanisms to identify salient features and relationships across the image set.
+This aligns with research on human attention, which suggests that holistic scene processing relies on rapid perception and attentional guidance Biederman (1987); Li et al.
+(2002).
+These mechanisms allow VLMs to selectively focus on relevant image features, mirroring human gaze patterns and attentional shifts.
+Similarly, the Deductive Rule Learning paradigm relies on the model’s ability to “memorize” the extracted rule for later application to query images, a process that likely involves both working memory Baddeley (2012) and long-term memory storage Squire (1992).
+Memory mechanisms are essential for encoding visual information, maintaining it during rule formulation, and retrieving it for comparison.
+This is reflected in the VLMs’ ability to store learned rules (long-term memory) and maintain representations of the images during analysis (working memory).
+Although we did not directly manipulate or measure these cognitive processes, their involvement is undeniable, and future research could explore these aspects more explicitly, perhaps by analyzing attention maps Bahdanau et al.
+(2016); Vaswani et al.
+(2017) or investigating the memory representations learned by VLMs during rule extraction.
+The interplay between attention and memory, crucial for both human and artificial intelligence, provides valuable insights into the computational underpinnings of visual reasoning Vaishnav and Serre (2023).
+Appendix A.3 Dataset
+A.3.1 Bongard OpenWorld Dataset
+We utilize a subset of 500 test cases from the Bongard OpenWorld dataset Wu et al.
 (2024), a benchmark for evaluating few-shot reasoning in machine vision.
-Each case presents a set of positive and negative examples, distinguished by an underlying "commonsense" rule.
-These rules, represented by numerical values (0-9), encode different visual and spatial concepts, as detailed below.
-B.2 Commonsense Value Categories
+Each of the 1001 Bongard samples in this dataset contains 7 positive and 7 negative images, distinguished by an underlying "commonsense" rule represented by numerical values (0-9).
+To create our evaluation set, we followed a specific sampling strategy.
+Each test case in our subset consists of 6 positive images, 6 negative images, and 1 query image.
+From each of the original 1001 Bongard samples, we can generate up to 14 unique, balanced test cases using combinations.
+For our evaluation, we selected the first 250 Bongard samples from the full dataset.
+From each of these 250 samples, we selected one positive and one negative test case, resulting in a total of 500 test cases for our analysis.
+The specific samples used in our analysis will be made publicly available upon acceptance of this work.
+These rules encode different visual and spatial concepts, as detailed below.
+Figure A.1: Illustration of a Bongard OpenWorld dataset showing three positive and negative samples along with a query image.
+Distinguishing rule in this example is A group photo at a wedding reception.
+Based on the rule, query image belongs to negative category.
+A.3.2 Commonsense Value Categories
 The Bongard OpenWorld dataset categorizes the underlying rules into several visual concept categories, some of which are explicitly related to commonsense reasoning.
@@
 Table A.1: Commonsense ID Categories in Bongard OpenWorld
-B.3 Commonsense Value Distribution in Our Subset
+A.3.3 Commonsense Value Distribution in Our Subset
 Table A.2 provides a detailed breakdown of the distribution of commonsense values within our 500 test-cases subset.
-As observed, the value ’0’ ("Anything else," marked as commonsense) is significantly more prevalent than other values.
-This distribution should be considered when analyzing the model’s performance, as it may indicate a bias towards learning certain types of rules, particularly those related to the "Anything else" category.
+As observed, the value ‘0’ (‘Anything else’, marked as commonsense) is significantly more prevalent than other values.
+This distribution should be considered when analyzing the model’s performance, as it may indicate a bias towards learning certain types of rules, particularly those related to the ‘Anything else’ category.
 ID Count Percentage
@@
 Total 500 100%
-Table A.4: Distribution of Commonsense ID in Bongard OpenWorld Dataset
-B.4 Dataset Availability
+Table A.2: Distribution of Commonsense ID in Bongard OpenWorld Dataset test cases used for experiments
+A.3.4 Dataset Availability
 The complete Bongard OpenWorld dataset is publicly available at Bongard-OpenWorld.github.io.
-Appendix C Evaluation Metrics
-Classification Metric:
-For evaluating the performance of the VLMs, we employ categorization accuracy on the binary classification task.
-This metric measures the percentage of test instances correctly classified as either positive or negative as present in the Bongard Problems.
-Semantic Similarity:
-We leverage a Semantic Similarity Evaluator to quantify the semantic relatedness between generated responses (e.g., question answers) and ground truth answers.
-This evaluator is inspired by the work of Risch et al.
-(2021), which explores semantic answer similarity for evaluating question answering models.
-Appendix D Model Details
+Appendix A.4 Model Details
 This section provides comprehensive details on the Vision-Language Models (VLMs) utilized in our experiments.
-D.1 Vision-Language Models (VLMs)
+A.4.1 Vision-Language Models (VLMs)
   • ChatGPT (GPT-4o with vision): OpenAI’s GPT-4o, accessed with vision capabilities via the openai library, processes both images and text seamlessly.
@@
 However, its closed-source nature and cost make it less accessible for extensive experimentation.
-  • Gemini (2.0-flash-exp): Google’s Gemini is a multimodal model Google (2024) emphasizing speed and computational efficiency in its "flash" version, accessed through the google.generativeai library.
+Accessed version of the model is 2024-08-06.
+  • Gemini (2.0-flash-exp): Google’s Gemini is a multimodal model Google (2024) emphasizing speed and computational efficiency in its “flash” version, accessed through the google.generativeai library.
 While it provides rapid inference capabilities, it is an experimental model with limited public documentation, restricting detailed architectural insights.
@@
 However, its relative novelty means limited community feedback and tooling support.
-D.2 Model Configurations and Hyperparameters
+A.4.2 Model Configurations and Hyperparameters
 For closed-source models, we utilized API calls by installing the openai, mistralai (for Pixtral), and google.generativeai libraries, enabling access to ChatGPT (GPT-4o with vision) OpenAI (2024), Pixtral (12B) Agrawal et al.
-(2024), and Gemini (2.0-flash-exp) Google (2024), respectively.
+(2024), and Gemini 2.0 (flash-exp) Google (2024), respectively.
 The open-source models, on the other hand, were accessed using the ollama Ollama (2025) library.
   • Input Format: Bongard problems were presented by encoding images in base64 and embedding them within text prompts.
-For instance: "Here are two sets of images.
+For instance: “Here are two sets of images.
 The first set represents positive examples, and the second set represents negative examples of a concept.
 What is the rule that distinguishes the two sets?
-"
-  • Image Size: For OpenAI and Gemini models, the default image sizes provided by their APIs were used. For all other models, the maximum image size was restricted to 1024 pixels to ensure uniformity.
+”
+  • Image Size: For OpenAI and Gemini 2.0 models, the default image sizes provided by their APIs were used.
+For all other models, the maximum image size was restricted to 1024 pixels to ensure uniformity.
   • Decoding Parameters: The temperature parameter was set to 0 to maintain deterministic outputs during evaluation.
@@
   • Hardware: Experiments were conducted using NVIDIA GPUs, including 2080Ti, 3090, and 6000 Ada, ensuring adequate computational power for all model evaluations.
-Appendix E Results
-Concept Category ID GPT 4.0 Gemini 2.0
-Anything else 0 92.88 94.23
-Human-Object Interaction (HOI) 1 86.67 92.86
-Taste / Nutrition / Food 2 100.00 85.71
-Color / Material / Shape 3 88.89 91.67
-Functionality / Status / Affordance 4 100.00 100.00
-And / Or / Not 5 90.00 80.00
-Factual Knowledge 6 90.00 90.00
-Meta Class 7 100.00 100.00
-Relationship 8 100.00 100.00
-Unusual Observations 9 92.31 92.31
-Table A.5: Accuracy per Commonsense ID and Categories in Bongard OpenWorld
-Gemini 2.0 GPT-4o
-Category Mean Std Mean Std
-Positive 0.915 0.02 0.9022 0.02
-Negative 0.868 0.02 0.8659 0.02
-Table A.6: Semantic Similarity between query image Descriptions and Identified Rules
-SL NO Task ID Caption Reason
-1 0021_neg_0 Cars on the city streets at night The image is synthetic, not a real-world photograph.
-2 0014_neg_0 A person playing a guitar. The image is synthetic, lacking real-world context.
-3 0033_neg_0 A bicycle is placed in the corner The image is synthetic and does not represent real-world scenarios.
-4 0037_neg_0 The girl has long and thin braids on her head. While technically correct, the braids are incorrectly interpreted or misrepresented.
-5 0076_pos_0 Various kinds of rings The classification error arises due to confusion between antique and artistic presentation styles.
-6 0076_neg_0 Various kinds of rings The classification error arises due to confusion between antique and artistic presentation styles.
-7 0082_neg_0 Live coral on the sea floor. The analysis is too fine-grained.
-8 0084_neg_0 A wooden fence surrounding a grassy field. Failure to accurately identify the wooden fence in the image.
-9 0112_neg_0 A wooden floor in the living room. The analysis is too fine-grained, and the image does not match the caption.
-10 0117_neg_0 Colorful ribbons. While ribbons are present, they are not colorful as described.
-11 0122_neg_0 A satellite view of Earth. Unable to recognize or correctly interpret the satellite image.
-12 0136_pos_0 Spectator seats view in the stadium. The analysis is overly detailed, leading to a categorization error.
-13 0213_neg_0 Checkerboard pattern fabrics Only the checkerboard pattern was identified, missing other key aspects.
-14 0234_neg_0 A beautiful stone sculpture The image is misclassified as a stone obelisk monument rather than a sculpture.
-15 0247_pos_0 Small river filled with reeds Although the identification is correct, the error stems from overly fine-grained analysis.
-Table A.7: Task IDs and descriptions where GPT-4o and Gemini both incorrectly categorize.
-E.1 Holistic Analysis
+A.4.3 Evaluation Metrics
+Classification Metric:
+For evaluating the performance of the VLMs, we employ categorization accuracy on the binary classification task.
+This metric measures the percentage of test instances correctly classified as either positive or negative as present in the Bongard Problems.
+Semantic Similarity:
+We leverage a Semantic Similarity Evaluator to quantify the semantic relatedness between generated responses (e.g., question answers) and ground truth answers.
+This evaluator is inspired by the work of Risch et al.
+(2021), which explores semantic answer similarity for evaluating question answering models.
+Appendix A.5 Model Prompts
+A.5.1 Direct Visual Rule Learning
+The prompt used for the Direct Visual Rule Learning paradigm is designed to elicit a holistic analysis of the provided images, encouraging the model to identify a distinguishing rule and apply it to the query image.
+The prompt emphasizes the distinction between positive ($cat\_2$) and negative ($cat\_1$) examples and guides the model to provide a structured output containing its analysis, the identified rule, details about the query image, and the final classification.
 def visual_concept_test_prompt(m, n):
@@
 - **Conclusion**: (cat_1 or cat_2)
-E.2 Deductive Rule Learning
-First-Stage Prompt (Rule Extraction)
+A.5.2 Deductive Rule Learning
+The Deductive Rule Learning paradigm employs a two-stage prompting strategy.
+The first stage focuses on rule extraction from positive and negative examples, while the second stage applies the extracted rule to classify a query image.
+The prompts for each stage are detailed below.
+A.5.2.1 First-Stage Prompt (Rule Extraction)
+This prompt guides the model to identify and summarize a distinguishing rule based on provided positive and negative examples. It emphasizes conciseness in the rule summary.
 def visual_concept_prompt(m, n):
@@
 raise
-Second-Stage Prompt (Rule Application)
+A.5.2.2 Second-Stage Prompt (Rule Application)
+This prompt presents the previously extracted rule summary and a query image, prompting the model to classify the image based on the rule.
+It reinforces the Bongard problem context and requests a structured output.
 # Define the visual analysis prompt
@@
 - **Conclusion**: (cat_1 or cat_2)
-E.3 Componential Analysis
-First-Stage Prompt (Image Description Generation)
+A.5.3 Componential Analysis
+The Componential Analysis paradigm also uses a two-stage prompting strategy.
+The first stage generates detailed image descriptions, while the second stage derives a rule from these descriptions and applies it to a query image.
+The specific prompts for each stage are presented below.
+A.5.3.1 First-Stage Prompt (Image Description Generation)
+This prompt instructs the model to generate a comprehensive, hierarchical description of a given image in JSON format.
+It guides the model to cover various aspects of the image, from scene and objects to activities and contextual elements, facilitating detailed comparative analysis in the subsequent stage.
 # Define the visual analysis prompt
@@
 Ensure that the JSON output is clear, well-formatted, and free of unnecessary explanations. Omit the ‘‘‘json tags at the beginning and end of the page.
-Second-Stage Prompt (Rule Derivation Instruction)
-def system_eval_prompt():
-return """
-You are an expert in visual reasoning and analysis, working with complex datasets such as the Bongard dataset.
-Your task is to analyze descriptions of images from two categories (cat_2 and cat_1) along with a query image, provided in JSON format.
-The goal is to:
-1. Identify the **common characteristics or patterns** in the positive samples (cat_2) that distinctly separate them from the negative samples (cat_1).
-2. Derive a **clear and concise rule** that defines the positive samples.
-3. Apply this rule to the query image and categorize it as belonging to either cat_1 or cat_2.
-Ensure the output is clear, concise, and follows the provided format:
-- **Analysis**: Provide a brief analysis of the distinguishing characteristics or patterns.
-- **Rule**: State the rule that separates cat_2 from cat_1.
-- **Query Image**: Analyze the query image based on the derived rule.
-- **Conclusion**: Clearly state whether the query image belongs to cat_1 or cat_2.
-Respond in the format provided, and avoid unnecessary explanations or text.
-Third-Stage Prompt (Rule Application and Classification)
+A.5.3.2 Second-Stage Prompt (Rule Derivation Instruction)
+This prompt guides the model to analyze the JSON descriptions generated in the first stage, derive a distinguishing rule, and apply it to classify a query image.
+It emphasizes the use of the provided JSON format and requests a structured output.
 def user_eval_prompt(all_image_specs, m, n):
@@
 - Cat_1: Negative samples that may not follow any specific rule.
-The image descriptions for the positive samples, negative samples, and the query image are provided in JSON format.
+The image descriptions for the positive samples, negative samples, and the test image are provided in JSON format.
 Analyze the common patterns or characteristics in the cat_2 samples that distinguish them from cat_1 samples.
@@
 1. Derive the rule that defines the cat_2 samples.
-2. Apply this rule to categorize the query image.
+2. Apply this rule to categorize the test image.
 Here are the image descriptions:
@@
 {all_image_specs[m:m+n]}
-### Query Image:
+### Test Image:
 {all_image_specs[-1]}
@@
 - **Rule**: (The distinguishing rule here)
-- **Query Image**: (Query image details)
+- **Test Image**: (Test image details)
 - **Conclusion**: (cat_1 or cat_2)
+Appendix A.6 Results
+This appendix provides supplementary information and detailed results for the experiments described in the main paper.
+We present additional performance breakdowns and analyses to offer a more comprehensive view of the models’ capabilities on the Bongard OpenWorld dataset.
+A.6.1 Componential Analysis Results by Commonsense Category
+Table A.3 presents a detailed breakdown of the overall accuracy achieved by GPT-4o and Gemini 2.0 on the Bongard OpenWorld test set using the Componential Analysis paradigm.
+The results are categorized by Commonsense ID, providing insights into model performance across different types of commonsense reasoning challenges.
+This granular view allows for a more nuanced understanding of the models’ strengths and weaknesses in handling various aspects of commonsense knowledge.
+ID Concept Category GPT 4o Gemini 2.0
+0 Anything else 92.88 94.23
+1 Human-Object Interaction (HOI) 86.67 92.86
+2 Taste / Nutrition / Food 100.00 85.71
+3 Color / Material / Shape 88.89 91.67
+4 Functionality / Status / Affordance 100.00 100.00
+5 And / Or / Not 90.00 80.00
+6 Factual Knowledge 90.00 90.00
+7 Meta Class 100.00 100.00
+8 Relationship 100.00 100.00
+9 Unusual Observations 92.31 92.31
+Table A.3: Overall accuracy of GPT-4o and Gemini 2.0 on the Bongard OpenWorld test set using Componential Analysis, broken down by Commonsense ID.
+A.6.2 Semantic Similarity in Deductive Rule Learning
+Table A.4 presents the mean and standard deviation of the semantic similarity scores between the descriptions of the query images and the rules derived by the models during the Deductive Rule Learning experiments.
+These scores provide a measure of how well the extracted rules capture the semantic content of the images and the underlying Bongard problem.
+Analyzing these similarities helps to understand the relationship between the models’ ability to articulate rules and their overall performance on the task.
+Gemini 2.0 GPT-4o
+Category Mean Std Mean Std
+Positive 0.915 0.02 0.9022 0.02
+Negative 0.868 0.02 0.8659 0.02
+Table A.4: Semantic Similarity between query image descriptions and identified Rules derived from Deductive Rule Learning
+A.6.3 Impact of Chain-of-Thought Prompting
+This section investigates the influence of Chain-of-Thought (CoT) prompting on VLM performance in the Direct Visual Rule Learning paradigm.
+CoT prompting, which encourages models to generate intermediate reasoning steps, has demonstrated significant improvements in complex reasoning tasks Wei et al.
+(2022b).
+We explore whether this approach is similarly beneficial in the visual domain, particularly for Bongard problems that demand a combination of perception and reasoning.
+To assess the impact of CoT, we compare two prompting strategies: one that explicitly guides the model through a chain-of-thought by requesting a structured output (analysis, rule, query image description, and classification), and another that simply asks for the final classification.
+Our analysis reveals a striking difference in performance between the two strategies.
+Without CoT prompting, the accuracy on the Direct Visual Rule Learning task drops significantly, from 80% to 61.60%.
+This decline is particularly pronounced for negative samples, where the accuracy falls from 66.40% to 39.20%, suggesting that the model struggles to identify instances where the rule does not apply when not explicitly prompted for a detailed reasoning process.
+This finding underscores the crucial role of CoT in guiding the model’s reasoning and enabling it to effectively navigate the complexities of visual reasoning tasks.
+These results are summarized in Table A.5.
+Model Accuracy (%)
+Overall neg pos
+Without CoT 61.60 39.20 84.00
+With CoT 80 66.40 93.60
+Table A.5: Impact of CoT Prompting on Direct Visual Rule Learning when evaluated using GPT-4o model
+The substantial impact of CoT observed in our experiments highlights its importance as a technique for eliciting more robust and reliable visual reasoning in VLMs.
+This suggests that incorporating CoT-like mechanisms into the model architecture or training process could lead to significant improvements in performance on complex visual tasks.
+Furthermore, our findings suggest that future research should explore the application of CoT prompting to other visual reasoning tasks and investigate the optimal strategies for integrating CoT into different VLM architectures.
+A.6.4 Error Analysis Examples
+This section provides detailed examples of Bongard OpenWorld test cases where both GPT-4o and Gemini 2.0 made incorrect classifications using the Componential Analysis paradigm.
+These examples illustrate common challenges faced by the models, offering insights into their limitations in visual reasoning and commonsense understanding.
+The analysis focuses on cases where both models failed, highlighting shared weaknesses.
+The “Reason” column explains the likely cause of misclassification based on the GPT-4o output (longer analysis, short rule, and test image description).
+These captions are reproduced from Wu et al.
+(2024).
+Table A.6: Examples of Bongard OpenWorld Test Cases with Captions (Reproduced from Wu et al.
+(2024)) and Reasoning for Incorrect Classifications for images classified incorrectly by componential analysis by both by GPT-4o and Gemini 2.0.
+Human error analysis is based on the GPT-4o textual output: a longer analysis, short rule and the test image description.
+SL NO Test ID Caption Reason
+1 0021_neg_0 Cars on the city streets at night Weak reasoning (by similarity): a specific object in the rule (vehicles) is missing in the test image (impressionist painting), although deducible.
+2 0014_neg_0 A person playing a guitar. Rule misses specific object (guitar) while being too general. The object is present in longer analysis.
+3 0033_neg_0 A bicycle is placed in the corner Rule and test image analysis miss a specific property (in a corner / standing against a wall with ground visible).
+The test image is a collage of drawings on a wooden background.
+4 0037_neg_0 The girl has long and thin braids on her head. Rule and test image analysis misses a specific property (specifically long and thin braids), being too general.
+5 0076_pos_0 Various kinds of rings Rule misses a specific object (ring), being too general.
+Instead, rule focuses on a specific property (intricate design), which is missing in the test image and its description.
+6 0076_neg_0 Various kinds of rings Rule misses a specific object (ring), being too general.
+7 0082_neg_0 Live coral on the sea floor. Weak reasoning (by similarity): rule finds specific object (coral), but test image analysis does not.
+8 0084_neg_0 A wooden fence surrounding a grassy field.
+Rule misses a specific object (grass), instead requiring greenery (grass or trees).
+The test image is a collage of natural images with greenery, no grass or trees are directly visible.
+9 0112_neg_0 A wooden floor in the living room. Rule misses two specific objects (living room and a wooden floor), while focusing on just one object (wooden) and generalizing too much (indoor).
+10 0117_neg_0 Colorful ribbons. Rule misses a specific object (colorful ribbons), generalizing too much.
+11 0122_neg_0 A satellite view of Earth. Rule misses a specific viewpoint (top-down satellite view), using a too general (aerial view) instead.
+12 0136_pos_0 Spectator seats view in the stadium.
+Weak reasoning: rule says “often active sports or spectators”, none of which are present in the test image or description.
+This lack is used as grounds for misclassification.
+13 0213_neg_0 Checkerboard pattern fabrics Rule misses a specific object (fabric), the specific versions of which are present in analysis.
+14 0234_neg_0 A beautiful stone sculpture Rule misses the correct property (tall), using instead another (prominent) of an obelisk. Test image features a prominent, but not a tall obelisk.
+15 0247_pos_0 Small river filled with reeds Rule misses a specific object (reeds) present in analysis in a longer list, while prohibiting wrong objects (industrial presence), which are distantly present in test image.